### v2

Abstract
In stars that lie on the main sequence in the Hertzsprung Russel diagram, like our Sun, hydrogen is fused to helium in a number of nuclear reaction chains and series, such as the proton-proton chain and the carbon-nitrogen-oxygen cycles. Precisely determined thermonuclear rates of these reactions lie at the foundation of the standard solar model.
This review, the third decadal evaluation of the nuclear physics of hydrogen-burning stars, is motivated by the great advances made in recent years by solar neutrino observatories, putting experimental knowledge of the pp chain neutrino fluxes in the few-percent precision range. The basis of the review is a one-week community meeting held in July 2022 in Berkeley, California, and many subsequent digital meetings and exchanges.
The relevant reactions of solar and stellar hydrogen burning are reviewed here, from both theoretical and experimental perspectives. Recommendations for the state of the art of the astrophysical S-factor and its uncertainty are formulated for each of them.
Several other topics of paramount importance for the solar model are reviewed, as well: recent and future neutrino experiments, electron screening, radiative opacities, and current and upcoming experimental facilities. In addition to reaction-specific recommendations, also general recommendations are formed.
Contents
Introduction
The present review summarizes the state of our understanding, in the third decade of the 21st century, of the nuclear reactions and decays taking place in solar and stellar hydrogen burning. It also addresses related issues, including solar neutrino detection, radiative opacities, electron screening of nuclear reactions, and the status of the current facilities for measuring cross sections and opacities.
As was the case for two previous decadal reviews [1998RvMP...70.1265A, 2011RvMP...83..195A], the present review summarizes the progress made over the past decade in advancing the nuclear physics of main-sequence stars and makes recommendations for future work.
I.1 Purpose
Ray Davis’s measurements of the flux of solar neutrinos [1968PhRvL..20.1205D, 2003RvMP...75..985D] showed a significant deficit with respect to the predictions of the standard solar model (SSM) [1966PhRvL..17..398B]. This discrepancy became known as the "solar neutrino problem." An intense debate ensued about possible explanations for this discrepancy, summarized in the entertaining review by 1988Natur.334..487B. Suggestions included plausible flaws in the SSM that might produce a cooler solar core (e.g. 1981A&A....96....1S; 1996PhRvL..77.4286C), new particle physics such as neutrino oscillations (1969PhLB...28..493G; 1978PhRvD..17.2369W), changes in the nuclear physics governing He synthesis, such as an unidentified low-energy resonance in the 3He+3He reaction affecting extrapolations of laboratory cross sections to solar energies [1972Natur.238...24F, 1972PhLB...40..602F], or the presence of dark matter particles modifying energy transport in the solar core [1985ApJ...294..663S]. By the mid-1990s, with new results from Kamioka II/III (1989PhRvL..63...16H; 1996PhRvL..77.1683F) and the SAGE/GALLEX (1991PhRvL..67.3332A; 1992PhLB..285..376A) experiments revealed a pattern of neutrino fluxes inconsistent with the expected scaling of those fluxes with the solar core temperature. Attention increasingly turned to neutrino oscillations and other potential particle physics solutions. Concerns about the nuclear physics of the SSM also evolved, focusing more on the uncertainties that might inhibit extraction of any such “new physics.”
At that time, during a meeting on the Solar Neutrino Problem hosted by the Institute for Nuclear Theory (INT), University of Washington, Seattle, a suggestion was made to convene the community working on solar nuclear reactions, in order to reach consensus on the best current values of cross sections and their uncertainties. The INT hosted the proposed workshop in February, 1997, drawing representatives from almost every experimental group active in this area, as well as many of the theorists who were engaged in solar neutrino physics. The working groups that formed during this meeting worked over the following year to evaluate past work, determining the needed cross sections and their uncertainties. The results of this evaluation, which came to be known as Solar Fusion I or 1998RvMP...70.1265A [1998RvMP...70.1265A], became the standard for use in solar modeling over the following decade.
An update of 1998RvMP...70.1265A was launched with the January, 2009, workshop “Solar Fusion Cross Sections II”, hosted again by the INT in Seattle. Initial results from SuperKamiokande, the Sudbury Neutrino Observatory, and Borexino were then in hand, and the conversion of approximately two thirds of solar electron neutrinos into other flavors had been firmly established [2002PhRvL..89a1301A]. Consequently, the motivation for the study was to ensure that solar model predictions would be based on the most current nuclear physics, so that meaningful uncertainties could be placed on neutrino parameters derived from solar neutrino measurements, such as the mixing angle . An additional source of uncertainty had arisen at the time, with the advent of three-dimensional (3D) radiative-hydrodynamic models of the Sun’s atmosphere for determining element abundances from photo-absorption lines [2006NuPhA.777....1A, 2009ARA&A..47..481A]. This improved analysis, though, lowered the inferred metallicities, leading to tension with the Sun’s interior sound speed profile determined from helioseismology [2002RvMP...74.1073C]. The discrepancy between higher metallicity models that accurately reproduce the Sun’s interior sound speed and lower metallicity models based on the most current treatment of the photosphere was termed the “solar composition problem”. While the neutrino fluxes were not known to the precision needed to distinguish between the competing models, it had been shown that the solar core metallicity could be extracted directly from future CN solar neutrino measurements, if the precision of associated nuclear cross sections were improved [2008ApJ...687..678H]. Reflecting the impact of new experimental work and improved theory, the 2011RvMP...83..195A study of 2011RvMP...83..195A significantly revised the 1998RvMP...70.1265A cross section for the driving reaction of the CNO-I cycle, and updated key pp-chain cross sections.
The present SF III review began with a four-day workshop in July, 2022, “Solar Fusion Cross Sections III.” The meeting was hosted by Physics Frontier Center N3AS (Network for Neutrinos, Nuclear Astrophysics, and Symmetries)111https://n3as.berkeley.edu/, a consortium of institutions involved in multi-messenger astrophysics, whose central hub is the University of California, Berkeley. Sessions were held at the David Brower Center, Berkeley, and the Physics Department, UC Berkeley, with the assistance of N3AS staff. The workshop was organized by N3AS, by Solar Fusion’s long-term sponsor, the INT, and by the European ChETEC-INFRA Starting Community for Nuclear Astrophysics222https://www.chetec-infra.eu/.
The 47 workshop participants represented most of the leading experimental and theoretical research groups that are active in the field. Workshop participants were organized into nine working groups, who were charged with reviewing and evaluating work completed since 2011RvMP...83..195A, to produce updated recommendations. The structure of this review largely follows that of the working groups. The timing of SF III was driven in part by planned large-scale asteroseismic surveys of hydrogen-burning stars such as ESA’s PLATO333https://platomission.com/ mission [2014ExA....38..249R], and the expectation that associated data analysis will require large libraries of stellar models in which masses, ages, metallicities, and other parameters are varied. SF III will provide the best current nuclear astrophysics input for such modeling. This review summarizes that input, based on the recommendations of the nine SF III working groups.
I.2 Terminology used
The terminology used here follows recent practice, as summarized for example in the NACRE compilation [1999NuPhA.656....3A]. It is briefly described to aid the reader.
Nuclear reactions are designated as
where is the target nucleus, is the bombarding particle, and and are the reaction products, with distinguished as the heavier product. Frequently in nuclear astrophysics, the reaction cross section is re-expressed in terms of the S-factor , with S related to as described below, and with the mass numbers of entrance channel nuclei and .
The relation between the nuclear cross section and astrophysical S-factor is
with and the charge numbers of the target and projectile, the fine structure constant, the relative velocity of the interacting particles in units of , the reduced mass, and the center-of-mass energy. (In this review we use energy units of MeV unless otherwise stated.) This removes the sharp energy dependence from the cross section associated with s-wave Coulomb scattering off a point nucleus, thereby accounting for the leading penetration effects associated with the Coulomb barrier. The quantity is known as the Sommerfeld parameter. Consequently, unless there are resonances, varies gently, and at low energies can generally be expressed as a low-order polynomial in .
The S-factor in Eq. (2) is defined in terms of the nuclear (or “bare") cross section, while laboratory experiments are performed with targets consisting of neutral atoms. As electrons shield the nucleus, lowering the Coulomb barrier, their presence enhances the laboratory cross section. The S-factors derived from laboratory data must be corrected for this effect. As the energy is lowered, the screening correction becomes increasingly important. Thus particular care must be taken when extrapolating laboratory cross sections to lower energies, as Fig. 4 of 2011RvMP...83..195A illustrates. Unless otherwise noted, the S-factors discussed here are those for bare nuclei.
The rapid rise in the nuclear cross section with energy, reflecting the higher probability of penetrating the Coulomb barrier, competes with a rapid decrease in the probability of finding interacting particles with the requisite center-of-mass energy in the high-energy tail of the Maxwell-Boltzmann distribution. Consequently, the thermonuclear reaction rate for a given stellar temperature takes the form
where is proportional to the Sommerfeld parameter of Eq. (2) and is Boltzmann’s constant. The maximum of the integrand defines the Gamow peak – the most probable energy for interactions – while the range around the peak is frequently termed the Gamow window.
For the solar reactions of interest, the energy at the Gamow peak is well below the height of the Coulomb potential. The slowly varying can be expanded as a power series around , with the leading term being , and with corrections given by derivatives taken at ,
Most of the results presented in this review will be given in terms of and its derivatives.
For most of the nuclear reactions studied here, the notation is adopted, where and are the mass numbers of the projectile and target nuclei, respectively. A summary of SF III reactions and recommended astrophysical S-factors are presented in Table 1.
Finally, we provide a short list of abbreviations that are frequently used throughout several sections:
- ANC
asymptotic normalization coefficient.
- BBN
Big Bang nucleosynthesis.
- CME
common-mode error.
- EFT
effective field theory, and variants: chiral (EFT), pion-less (EFT), hybrid (EFT*), etc.
- LO, NLO, NNLO
leading order, next to (next to) leading order, etc.
- SSM
Standard Solar Model.
| Reaction | (MeV b) | (b) | (MeV-1 b) | Section | |
| III | |||||
| see text | IV | ||||
| V | |||||
| – | VI | ||||
| – | – | VII | |||
| – | – | IX | |||
| – | – | LABEL:subsec:WG3:Total_Sfactors | |||
| LABEL:sec:OtherCNO:CNO1:c12pg | |||||
| LABEL:sec:OtherCNO:CNO1:c13pg | |||||
| LABEL:sec:OtherCNO:CNO2:n15pg | |||||
| LABEL:sec:OtherCNO:CNO1:n15pa | |||||
| LABEL:sec:OtherCNO:CNO2:o16pg | |||||
| – | – | LABEL:sec:OtherCNO:CNO2:o17pg | |||
| – | – | LABEL:sec:OtherCNO:CNO3:o18pg | |||
| – | – | LABEL:sec:OtherCNO:NeNa:ne20pg | |||
| – | – | LABEL:sec:21nepg | |||
| – | – | LABEL:sec:OtherCNO:NeNa:ne22pg | |||
| LABEL:sec:OtherCNO:NeNa:na23pg |
I.3 Scope and structure of the review
In the Sun, the primary mechanism for the conversion of four protons into 4He is the pp chain of Fig. 1, during which two charge-changing weak interactions take place, each converting a proton into a neutron with the emission of a neutrino. The pp chain consists of three main branches, ppI, ppII, and ppIII. These branches are distinguished by their dependence on the solar core temperature and by the neutrinos they produce, with the pp, 7Be, and 8B neutrinos serving as “tags” for the ppI, ppII, and ppIII chains, respectively. The SSM predicts that the fluxes of pp, 7Be, and 8B neutrinos have an approximate relative temperature scaling of 1:T:T, where Tc is the solar core temperature [annurev:/content/journals/10.1146/annurev.aa.33.090195.002331]. By the mid-1990s, solar neutrino experiments established that the measured fluxes differed significantly from this expected pattern. This contributed to growing expectations that new neutrino physics might be the solution of the solar neutrino problem. It also was an important motivation for 1998RvMP...70.1265A, as errors in the nuclear physics of the pp chain could distort the relationship between the various fluxes.
The pp chain accounts for 99% of solar hydrogen burning and are the main focus of this review. The remaining 1% is generated through the CNO-I cycle, see Fig. 2). Motivated in part by Borexino’s recent success in measuring the flux of solar CN neutrinos, this review also considers in more detail than past Solar Fusion reviews the reactions driving hydrogen burning in stars more massive than our Sun.
The review is structured as follows. Section II begins with a summary of past solar neutrino experiments, highlighting some of the open questions that remain. It discusses the second-generation experiments Super-Kamiokande, the Sudbury Neutrino Observatory, and Borexino that contributed in the resolution of the solar neutrino problem. It also describes Hyper-Kamiokande, SNO+, JUNO, and DUNE, new detectors in various stages of construction.
The cross section of the driving reaction of the pp chain, , is too small to be measured and thus must be taken from theory. This important reaction is reviewed in section III. With the exception of the minor hep branch, all other reactions and decays have been studied in laboratories, with measurements then combined with theory to predict rates under solar conditions. The 2H(He, 3He(3He,2)4He, and 3HeB reactions are reviewed in Secs. IV, V, and VI, respectively. The hep reaction is reviewed in Section VII, and the electron-capture reactions on pp and 7Be are discussed in Section VIII. The 7Be(,B reaction responsible for the high energy neutrinos measured in Super-Kamiokande is treated in Section IX.
While a modest contributor to solar energy production, the CNO-I cycle is a potentially important probe of core metallicity as its hydrogen burning is catalyzed by the Sun’s primordial C and N (Section LABEL:sec:OtherCNO).
A current focus of nuclear astrophysics laboratories is the higher-temperature pathways for hydrogen burning that operate in massive stars. Many of the contributing reactions have been re-measured. That progress is reviewed in Section LABEL:sec:OtherCNO.
Charged-particle nuclear reactions occurring at energies below the Coulomb barrier are affected by electron screening, the shielding of the nuclear charge by electrons. Because of their very different atomic environments, laboratory reactions and those occurring in the solar plasma are affected in distinct ways. Section LABEL:sec:Screening describes the current status of efforts to account for these differences, when solar cross sections are extracted from laboratory measurements.
The energy transport inside the radiative zone is affected by the metals444In this work, we follow the astrophysical terminology where ”metals” refer to all chemical elements except for the two lightest ones, hydrogen and helium. that were incorporated into the Sun when it first formed. While metals comprise less than 2% of the Sun by mass, they play an outsized role in determining the opacity. Very few of the needed radiative opacities have been measured, and the conditions under which measurements are made typically are not identical to those in the Sun. Section LABEL:sec:Opacities describes the current state of the art, and discusses connections between opacities and solar composition that have complicated efforts to resolve the solar metallicity problem.
Much of the progress made in constraining the reactions of the pp chain and CNO-I cycles has been possible thanks to new laboratories and the facilities they host. The underground LUNA laboratory is a prominent example [2018PrPNP..98...55B]. Section LABEL:sec:Facilities describes the capabilities of selected laboratories involved in the measurements reviewed here.
Solar neutrino observations
Solar neutrinos offer a unique tool box for probing both the fundamental properties of these elusive particles, and their interactions with matter, as well as understanding their source: the fusion reactions that power our Sun. The original motivation for observations of solar neutrinos was precisely the hope to probe solar fusion. The first successful experiment beginning in 1967 at the Homestake mine [1998ApJ...496..505C] offered the surprising result of a neutrino flux suppressed to approximately one third of expectation. The intervening years have seen a tour de force of experimental efforts, leading to the resolution of the Solar Neutrino Problem and the confirmation of neutrino oscillations [1998PhRvL..81.1562F, 2001PhRvL..87g1301A, 2002PhRvL..89a1301A, 2003PhRvL..90b1802E, 2004PhRvL..92r1301A]. These high-precision flux and spectral measurements mapped out the details of solar fusion, probed the structure of the Sun, and demonstrated (together with atmospheric neutrino measurements) that neutrinos have masses and mix. Table 2 lists the pp-chain and CNO cycle neutrino sources, their endpoint ( decay) or line (electron capture) energies , and predicted SSM fluxes.
| Source | Eν (MeV) | GS98 | AAGS21 | MB22p | units (cm-2s-1) |
| pp | 0.420 | 5.96 (0.6%) | 6.00 (0.6%) | 5.95 (0.6%) | 1010 |
| pep | 1.442 | 1.43 (1.1%) | 1.45 (1.1%) | 1.42 (1.3%) | 108 |
| hep | 18.77 | 7.95 (31%) | 8.16 (31%) | 7.93 (30%) | 103 |
| 7Be | 0.862(a) 0.384(b) | 4.85 (7.4%) | 4.52 (7.3%) | 4.88 (8.1%) | 109 |
| 8B | 15 | 5.03 (13%) | 4.31 (13%) | 5.07 (15%) | 106 |
| 13N | 1.198 | 2.80 (16%) | 2.22 (13%) | 3.10 (15%) | 108 |
| 15O | 1.732 | 2.07 (18%) | 1.58 (16%) | 2.30 (18%) | 108 |
| 17F | 1.738 | 5.35 (20%) | 3.40 (16%) | 4.70 (17%) | 106 |
Notes: (a) 90% and (b) 10% of the 7Be neutrino flux, respectively.
II.1 Open questions
The solar neutrino spectrum has been measured for energies above 233 keV, with some fluxes determined to a precision equivalent to or better than theoretical predictions. However, improved measurements can offer further insight on several aspects. For instance, more precise determinations of the fluxes of neutrinos from the CNO cycle (Fig. 2) could determine the contemporary abundances of N and O in the solar core [2007A&A...463..755C, 2011ApJ...743...24S, 2022A&A...667L...2K]. While these abundances differ significantly from the primordial values that are input into the SSM –the primordial abundances are altered by out-of-equilibrium CNO-cycle burning that occurs in the early Sun [Roxburgh85]– this burning redistributes the C, N, and O, but does not alter the net number of such nuclei. Consequently, CNO neutrino flux measurements might allow one to experimentally address the solar composition problem discussed below.
Solar neutrino flux measurements can also constrain physical processes in stars such as chemical mixing. The fluxes are sensitive to solar/stellar model inputs, including the radiative opacities (see Sect. LABEL:sec:Opacities). Measurements of the neutrino fluxes from 8B and 7Be can offer a handle on the temperature of the solar core, and other environmental factors [1996PhRvD..53.4202B].
A measurement of the pp flux with percent-level precision would provide a test of the luminosity constraint, further probing solar power generation mechanisms [2002PhRvC..65b5801B, 2021JPhG...48a5201V]. The one branch of the pp chain that yet remains undetected, the hep flux, has both the highest energy but also the lowest flux of neutrinos from the Sun. The spectrum extends beyond that of the 8B neutrinos, offering a small window for unequivocal observation. Limits have been placed on this flux [2020PhRvD.102f2006A], but a definitive observation will likely require a large, next-generation detector such as Hyper-Kamiokande (HyperK) [2018arXiv180504163H] or the Deep Underground Neutrino Experiment (DUNE) [2019PhRvL.123m1803C].
Finally, the observation of monochromatic neutrinos produced by electron capture reactions on 13N, 15O and 17F (which we refer to as ecCNO neutrinos) [1990PhRvD..41.2964B, 2004PhRvC..69a5801S, 2015PhLB..742..279V] could be rewarding in terms of physical implications. Indeed, they can be used as probes of the metallicity of the solar core. Moreover, they provide a measure of the electron neutrino survival probability at specific neutrino energies MeV within the broad transition region between vacuum and matter-enhanced oscillations, where currently we have no constraints. This would constitute a new test of the large-mixing-angle (LMA) MSW flavor oscillation paradigm. The detection of this subdominant component of the CNO-cycle is extremely difficult but it could be within reach of future very large ultra-pure liquid scintillator detectors [2015PhLB..742..279V].
Solar neutrinos provide an opportunity to understand the interaction of neutrinos with matter. The effect of MSW oscillations has a significant impact on the observed solar neutrino spectrum: with vacuum oscillation dominating at low energy, below approximately 1 MeV, where the survival probability is roughly one half, while matter effects further suppress the flux to a survival probability of roughly one third above approximately 5 MeV. The transition region between these two regimes offers an extremely sensitive probe of the details of the interactions of neutrinos with matter, including the potential to search for new physics such as sterile neutrinos, or non-standard interactions, by looking for distortions to the expected spectral shape [2019arXiv190700991B, 2020JHEP...02..038B, 2023JHEP...08..032C].
The same theory of matter effects predicts a small regeneration of electron neutrinos during the night time as they propagate through the bulk of the earth. This so-called “day/night effect” has been sought after by both Super-Kamiokande (SuperK) and the Sudbury Neutrino Observatory (SNO) [2014PhRvL.112i1805R, 2002PhRvL..89a1302A], but a significant observation is still limited by statistics. Future data from HyperKamiokande or DUNE may be needed to confirm our understanding of this prediction of the MSW effect.
The slight tension between measurements of the mass splitting parameters, in solar neutrino experiments and terrestrial data sensitive to the same parameter, from the KamLAND reactor experiment has now disappeared with the addition of new data on day-night effects and spectral shapes [Esteban2020TheFO, 2018PrPNP.102...48C]. Further tightening of the oscillation parameters can be expected from JUNO, which will obtain highly precise data on both the solar mass splitting and the mixing angle [2016JPhG...43c0401A].
II.1.1 The solar composition problem
Solar photospheric abundances, determined with spectroscopic techniques, are a fundamental input for the construction of SSMs, but also they are used as input to nearly all models in astrophysics, including stellar evolution, protoplanetary disks, and galactic chemical evolution. Over the last two decades, the development of three dimensional radiation hydrodynamic models of the solar atmosphere [2009MmSAI..80..711L, 2000A&A...359..729A] and of techniques to study line formation under non-local thermodynamic conditions [2005ARA&A..43..481A], together with improved descriptions of atomic properties (e.g. transition strengths 2022A&A...661A.140M, 2021MNRAS.508.2236B), have led to a significant revision of solar abundances. Initial results based on these more sophisticated methods favored a markedly lower solar metallicity [2009ARA&A..47..481A], particularly in the CNO elements, than obtained in the 90s from older techniques [1993oee..conf...15G, 1998SSRv...85..161G].
Recently, two groups have revisited the solar photospheric abundances using modern methods [2021A&A...653A.141A, 2022A&A...661A.140M]. While 2021A&A...653A.141A obtain results consistent with their previous findings, 2022A&A...661A.140M find an O abundance intermediate between that of Asplund’s group and those from the 90s, in agreement with another 3D-based determination by 2011SoPh..268..255C. Interestingly, 2022A&A...661A.140M find higher C and N abundances and, indirectly, a higher Ne abundance due to the larger Ne to O ratio measured in the solar corona [2018ApJ...855...15Y]. This leads to a combined metal-to-hydrogen ratio that is by chance comparable to those from 1993oee..conf...15G and 1998SSRv...85..161G, albeit with a different mixture of elements.
Considering that uncertainties in element abundances are difficult to quantify, it has become customary to consider two canonical sets of abundances, which we refer to as high metallicity (HZ) and low metallicity (LZ) solar admixtures. See, for example, 2017ApJ...835..202V for SSM reference values. In this context, the new solar abundance determinations by 2021A&A...653A.141A fall into the LZ category, while those from 2022A&A...661A.140M are HZ. Solar models employing the LZ abundances fail to reproduce most helioseismic probes of solar properties. This disagreement constitutes the so-called solar composition problem [2004ApJ...606L..85B, 2005ApJ...618.1049B, 2006ApJ...649..529D] that has defied a complete solution. All proposed modifications to physical processes in SSMs offer, at best, only partial improvements in some helioseismic probes, see e.g. [2005ApJ...627.1049G, 2007A&A...463..755C, 2008PhR...457..217B, 2010ApJ...713.1108G, 2011ApJ...743...24S]. The same conclusions are obtained with SSMs computed with the newest LZ [2021A&A...653A.141A] and HZ [2022A&A...661A.140M] abundances as discussed also there. An alternative possibility is to consider modifications to the physical inputs of SSMs at the level of the constitutive physics, radiative opacities in particular. This is possible because most helioseismic probes depend actually not directly on the solar composition, but on the radiative opacity profile in the solar interior, i.e. on the combination of solar composition and atomic opacities (see § LABEL:sec:Opacities). The same can be said for solar neutrinos from the pp chain. Early work [2005ApJ...621L..85B, 2004ESASP.559..574M] already suggested that a localized increase in opacities could solve or, at least, alleviate the disagreement of low-Z solar models with helioseismology, and 2009A&A...494..205C and 2010ApJ...724...98V showed that a tilted increase in radiative opacities, with a few percent increase in the solar core and a larger (15-20%) increase at the base of the convective envelope could lead to LZ SSMs that would satisfy helioseismic probes equally well as HZ SSMs.
The degeneracy between solar composition and opacities can be broken using CNO solar neutrinos, e.g., following the methodology developed in 2008ApJ...687..678H. Such a study was recently carried out by the Borexino collaboration [2020Natur.587..577B, 2022PhRvL.129y2701A, 2023PhRvD.108j2005B]. The Borexino measurement of the CNO-I cycle neutrino flux (the 13N and 15O fluxes) were used to determine the C+N core abundance. Results show a tension with LZ metallicity determinations, while being in better agreement with HZ mixtures. While the error budget is presently dominated by the uncertainty of the Borexino CNO neutrino measurement, a significant contributor to the error () is nuclear, due to uncertainties in , , and . The interpretation of future improved CNO neutrino measurements will be impacted, unless these nuclear physics uncertainties are reduced.
Very recently, 2024JHEP...02..064G presented a new global determination of all solar neutrino fluxes using all available experimental data, including the latest phases of Borexino. Results from this global analysis are in line with those from Borexino, although the added 13N and 15O fluxes are about 10% lower than the Borexino result alone. A comparison of the solar fluxes with SSM calculations [B23Fluxes] shows that HZ SSMs are in better agreement with solar neutrino fluxes than the LZ SSMs, pointing toward a C+N solar core abundance consistent with HZ abundances. The discrimination that solar neutrino fluxes can offer between solar compositions is at most, however, of the order of 2. (See Table 2 in 2024JHEP...02..064G).
Helioseismology also offers the potential to determine the total solar metallicity, i.e. without disentangling individual element abundances, rather independently from opacities. This relies on using the so-called adiabatic index , which deviates from 5/3 in regions of partial ionization. The underlying technique has been used widely to determine the helium abundance in the solar envelope and it has also been extended to determine solar metallicity. Attempting the latter is difficult because the imprint of partial ionization of metals is quite subtle. Previous work along this line [2006ApJ...644.1292A] found an overall metallicity consistent with HZ abundances, but new work [2024A&A...681A..57B] claims to favor LZ values. While this method depends very weakly on radiative opacities, the abundance determination is degenerate with the equation of state. Independent confirmation of these results would be desirable. In short, the controversy related to the solar composition is far from being resolved, currently with different indicators showing contradictory results.
II.1.2 The Gallium anomaly
With the aid of very intense radioactive sources of 51Cr and 37Ar, tests have been made of the rate of production of 71Ge by neutrino interactions on 71Ga, the basis of radiochemical measurements of the low-energy solar neutrino flux. Initial experiments showed lower rates than expected, with interesting but inconclusive statistical precision. Very recently the BEST experiment [2022PhRvL.128w2501B, 2022PhRvC.105f5502B, 2018PhRvD..97g3001B] has confirmed this anomaly at more than 4. Many possible explanations have been explored, but for each there are contradictions [2023EPJC...83..578K, 2022JHEP...11..082A, 2023PhLB..84237983G, 2023JHEP...05..143B]. A sterile neutrino explanation is disfavored [2022JHEP...10..164G, 2022EPJC...82..116G], particularly because of conflict with solar neutrino limits (see Fig. 3), as well as cosmological bounds [2021PhRvD.104l3524H].
The sterile neutrino contribution to the solar flux is fundamentally limited by the luminosity constraint, that the Sun’s total energy output is the result of nuclear reactions that produce neutrinos, whether active or sterile. If precise neutral-current data were available across the solar spectrum, a completely model-independent limit on a possible sterile component could be determined. The present neutral-current data at low energies (from Borexino) are imprecise but can be supplemented by much more precise charged-current data together with 3-flavor oscillation physics. Still better limits, at the cost of some model dependence, can be obtained with fits of experimental data to solar models, as shown in Fig. 3 above, as well as Fig. 1 of 2022EPJC...82..116G.
The precisely known rate of electron capture on 71Ge to 71Ga, for which the half-life is 11.43(3) d [1985PhRvC..31..666H], places an important constraint on the neutrino absorption rate [2023PhLB..84237983G, PhysRevC.108.035502]. The half-life has now been remeasured by two groups, 2024PhRvC.109e5501N and PhysRevC.108.L021602, yielding values of 11.468(8) d and 11.46(4) d, consistent with the previous determination while increasing the precision. The allowed matrix element that governs neutrino capture to the ground state of 71Ge can be extracted to a precision of 1% from the electron capture rate. The ground-state transition by itself generates a significant anomaly: inclusion of the excited-state contributions doubles the effect. In Appendix A we describe some of the details of both the 1997 extraction of the allowed (Gamow Teller) strength by 1997PhRvC..56.3391B and the recent update of PhysRevC.108.035502, ELLIOTT2024104082.
PhysRevC.108.035502, ELLIOTT2024104082 find the ground-state cross section for absorbing 51Cr neutrinos is 2.5% lower than Bahcall’s value. This work includes the contributions of weak magnetism and radiative corrections, which are shown to be sub-1% effects. An improved extraction of the excited-state contribution to neutrino absorption was performed using data from forward-angle (p,n) scattering. This led to a slight increase in that contribution relative to 1997PhRvC..56.3391B, thereby reducing the 2.5% difference above by about half, for the total 51Cr and 37Ar capture cross sections.
The procedures followed by PhysRevC.108.035502, ELLIOTT2024104082 could be extended to the solar neutrino 71Ga capture cross section, including the contribution from the high-energy 8B neutrinos. That has not been done, but as noted in Appendix A, could help resolve small discrepancies noted there.
Results from the SAGE and GALLEX/GNO experiments remain part of the solar neutrino database used in various global fits to neutrino parameters. A resolution of the anomaly that emerged from BEST and the four gallium detector calibration experiments is important as it would increase confidence in these data.
II.1.3 The Boron-8 neutrino spectrum
The neutrino spectrum from , extending to approximately 17 MeV, has been the most accessible part of the solar neutrino spectrum [2016PhRvD..94e2010A, 2020PhRvD.102f2006A], playing an important role in disclosing the physics of neutrino oscillation and testing the SSM. Knowing it with precision, particularly at the high energy end, has renewed interest in the context of observing the hep neutrinos [2006ApJ...653.1545A, 2020PhRvD.102f2006A], which could be accessible with future observatories [2022arXiv220212839A].
In 2011RvMP...83..195A it was decided to recommend the spectrum calculated by Winter et al. [2003PhRvL..91y2501W, 2006PhRvC..73b5503W, Win:2007] based on their measurement of the alpha spectrum from . This spectrum showed excellent agreement with an independent experiment [2006PhRvC..73e5802B]. Both radiative and recoil-order corrections were included by Winter et al. The radiative corrections are relatively small, due to a cancellation between the real and virtual contributions [1995PhRvD..52.5362B]. The recoil-order corrections are dominated by the weak magnetism part, which has been deduced from measurements of the analog electromagnetic decays [1995PhRvC..51.2778D].
Since then three additional measurements [2011PhRvC..83f5802K, 2012PhRvL.108p2502R, 2023PhRvC.107c2801L] of the -decay alpha spectrum have been performed. All three find that the peak of the alpha spectrum appears about 20 keV lower than determined by Winter et al. Overall, these new measurements yield differences in the neutrino spectrum of 5% below MeV, the energy above which the contribution becomes small compared to the hep component. Figure 4 shows the 8B spectral uncertainties on the scale of the hep spectrum. With resolution effects considered, it will be difficult to extract the hep flux from solar neutrino measurements without further reduction in 8B uncertainties.
In producing a recommendation for the neutrino spectrum we considered the following:
1.
Apparently the uncertainty estimations for the neutrino spectrum in 2012PhRvL.108p2502R are not quite correct. 2023PhRvC.107c2801L show that the uncertainty due only to the weak magnetism part, for which all authors follow similar prescriptions, is larger in the 0-12 MeV range than the overall uncertainty estimated by Roger et al. We estimated the uncertainties due to the weak magnetism term only and found agreement with those of Longfellow et al.
2.
All three of the recent efforts [2011PhRvC..83f5802K, 2012PhRvL.108p2502R, 2023PhRvC.107c2801L] used detectors (Si strip) more complicated than those used by Winter et al. (a Si surface barrier detector). The strip detectors allow for better position resolution and reduced beta summing. However, for detection from sources external to the detectors (as opposed to detection from sources implanted deep into a detector), the complicated pattern of dead layers needs to be measured and taken into consideration for accurate calibrations. Effects of partial charge collection in-between strips are also significant. The effects are different for a calibration source than from the emission from trapped as an ion.
Based on these considerations, we recommend the spectrum of 2023PhRvC.107c2801L, but suggest that conclusions sensitive to the choice of neutrino spectrum, such as evidence for hep neutrinos, should be evaluated using both the Winter et al. and the Longfellow et al. spectra.
Given the importance of the weak magnetism contributions, it would be good to have a new experiment with reduced uncertainties that could be compared directly to the results of 1995PhRvC..51.2778D. Recoil-order effects in the system were most recently measured by 2011PhRvC..83f5501S, making use of the alignment- correlation. Using combinations of their measurements and correlation data from 1980PhRvC..22..738M, they were able to extract some of the recoil-order matrix elements. Calculations of these matrix elements in the symmetry-adapted no-core shell model were recently performed by 2022PhRvL.128t2503S. The comparison to the experimentally derived matrix elements of 1980PhRvC..22..738M, 1995PhRvC..51.2778D, and 2011PhRvC..83f5501S is not straightforward because the experiments quote averages over the whole beta spectrum as opposed to fits for each of the four levels used in the calculation. A more detailed comparison of theory to experiment that takes into account the averaging would be helpful. An additional point of comparison that has been used previously is the beta spectrum from [1996PhRvC..54..411B]. An experiment with improved statistics and systematics that extends over the full range of beta energies would be very useful. The existing data from 1987PhRvC..36..298N is in good agreement with the alpha spectrum from Winter et al. But given the apparent 20 keV difference that the newer alpha-spectrum experiments described above report, it would be good to have a modern beta spectrum measurement specifically designed to address this difference.
II.2 Experimental program on solar neutrinos
A broad range of technology can be used to interrogate solar neutrinos. Radiochemical experiments utilizing chlorine and gallium played a crucial role in the early days of the solar neutrino problem. Large monolithic detectors have achieved great success in real-time measurements – from the water Cherenkov detectors such as Kamiokande, SNO, and SuperK, which could use the unique topology of Cherenkov light to point back to the neutrinos’ origin, to liquid scintillator detectors such as KamLAND and Borexino, whose high light yield and low threshold allow for precision spectroscopy.
II.2.1 Super-Kamiokande and Hyper-Kamiokande
The first real-time detection of solar neutrinos was achieved by the Kamiokande experiment [1989PhRvL..63...16H], which detected the neutrinos via elastic scattering of electrons, a process in which the outgoing electron’s direction is highly correlated to that of the incoming neutrino. This allowed Kamiokande to directly point back to the neutrinos’ origin: our Sun. The successor experiment, Super-Kamiokande, is a 50-kton volume of pure water, surrounded by over 10,000 photon detectors, which has achieved an unparalleled program of neutrino observations and other physics over its several decades of operation. This program has included the highest-precision measurement of the elastic scattering signal from solar neutrino interactions, as well as sensitive searches for the day/night distortions of the spectral shape [2016PhRvD..94e2010A]. Now filled with a gadolinium additive to enhance neutron capture, SuperK’s primary current focus is the search for the Diffuse Supernova Neutrino Background (DSNB) via inverse beta decay [2023ApJ...951L..27H]. It also continues its atmospheric neutrino program: SuperK measurements made 25 years ago demonstrated the neutrino oscillations were responsible for the puzzling zenith-angle-dependence of this flux [PhysRevLett.81.1562].
In parallel to ongoing operation of SuperK, an even larger sister project is under construction. At over 250 kton in total mass, and with improved light collection relative to SuperK, Hyper-Kamiokande (HyperK) will impact statistics-limited searches the involve the higher energy 8B solar neutrinos, including day/night effects, the search for hep neutrinos, and the shape of the 8B neutrino spectrum. After 10 years of operations, HyperK will reach a sensitivity to day/night effects 4 (8) , given the oscillation parameters deduced from reactor (solar) experiments. HyperK will serve as the far detector for JPARC’s long-baseline neutrino program.
II.2.2 Sudbury Neutrino Observatory and SNO+
SNO’s unique use of heavy water as a target medium offered the ability to detect solar neutrinos via two additional interactions besides elastic scattering. In the charged current interaction an electron neutrino interacts with the neutron in the deuteron, producing an electron and a proton. This interaction path is sensitive only to electron-flavor neutrinos at the few-MeV energy scale of solar neutrinos, providing a measurement of the pure flux. SNO also had access to the neutral current interaction, in which a neutrino of any flavor interacts with the deuteron, breaking it apart into its constituent nucleons. Being equally sensitive to all active flavors, the neutral current interaction yields a measurement of the total, flavor-blind neutrino flux. It was this capability that allowed SNO to resolve the solar neutrino problem, demonstrating that most of the s produced by the Sun were arriving at earth in a different flavor state [2001PhRvL..87g1301A, 2002PhRvL..89a1301A, 2004PhRvL..92r1301A].
SNO ceased data taking in 2006. Analysis of the data set has continued, and produced a number of new results, including constraints on non-standard effects such as Lorentz violation and neutrino decay [2018PhRvD..98k2013A, 2019PhRvD..99c2013A], limits on the fluxes of hep neutrinos and the diffuse supernova neutrino background (DSNB) [2020PhRvD.102f2006A], and neutron production from cosmogenic muons and from atmospheric neutrino interactions [2019PhRvD..99k2007A, 2019PhRvD.100k2005A].
After decommissioning, the detector was repurposed for the SNO+ experiment, in which the target material was replaced with a pure organic liquid scintillator: linear alkyl benzene (LAB), loaded with 2.2 g/L of the secondary fluor, PPO. The high light yield of this scintillator, coupled with the location in SNOLAB –at 6-km water equivalent, one of the deepest underground laboratories in the world– and an extremely well-understood detector, allow for a range of high-precision measurement programs. In a preliminary water phase, SNO+ demonstrated detection of neutron captures on hydrogen [PhysRevC.102.014002], an impressive technical achievement in an unloaded water detector; a low background measurement of 8B solar neutrinos [PhysRevD.99.012012]; several searches for invisible modes of nucleon decay [PhysRevD.99.032008, PhysRevD.105.112012]; and detection of antineutrinos from reactors over 200 km away [PhysRevLett.130.091801]. The low backgrounds achieved and the detector’s demonstrated technical capabilities set the stage for future measurements. In a “partial-fill” stage, in which the upper half of the detector was filled with liquid scintillator while the lower half still contained water, the first demonstration of event-level direction reconstruction was achieved for 8B neutrinos in a scintillator detector [PhysRevD.109.072002]. This was facilitated by a lower loading of PPO at that time, which results in lower scintillation yield and a slower time profile, effectively enhancing the Cherenkov component, which can be leveraged for directional information.
Now fully filled with liquid scintillator, SNO+ is the deepest, largest operating liquid scintillator detector in the world. The future program will include measurements of several solar neutrino fluxes, as well as the 8B neutrino energy spectrum.
II.2.3 Borexino
Water Cherenkov experiments such as Kamiokande, SuperK and SNO provided the first real-time measurement of solar neutrinos but their relatively high energy threshold made them sensitive only to a small fraction of the solar neutrino flux. To study in real time the bulk of solar neutrino emission a different detector technology was required. The Borexino experiment, located deep underground at the INFN Laboratori Nazionali del Gran Sasso, used an organic liquid scintillator target made of pseudocumene with PPO to detect the elastic scattering of solar neutrinos off electrons. The high light yield of the liquid scintillator made it possible to significantly lower the energy threshold, but given the small signal rate and the lack of a clear signature to separate it from the background (such as the direction indicated by the emission of Cherenkov radiation), the experiment required a long preparatory phase to develop the most advanced techniques to suppress the background, especially the one due to radioactive contamination of the liquid scintillator itself.
The detector’s extreme radiopurity was key to its success, and over its 14 years of data taking, ending in October 2021, Borexino proved itself capable of covering the entire solar neutrino spectrum. At the beginning of its data taking in 2007 the level of 238U and 232Th contamination in Borexino were lower than and respectively [2008PhLB..658..101B], paving the way for the first measurement of the 7Be sub-MeV solar neutrinos, followed by a low-threshold measurement of the 8B flux and a first indication of the pep neutrinos. After a purification campaign that further reduced the liquid scintillator contamination, Borexino aimed at improving the accuracy of its results in the measurement of the pp-chain solar neutrinos. The flux of neutrinos produced in the pp fusion process was first reported in 2014Natur.512..383B and later improved, along with all the other fluxes from the pp chain, apart from the hep neutrinos [2018Natur.562..505B, 2019PhRvD.100h2004A, 2020PhRvD.101f2001A]. The pp flux was determined to an uncertainty of 11%, improving the neutrino-based estimate of solar luminosity, while 7Be neutrinos were measured with an uncertainty of , half that of the SSM prediction. Furthermore, using the SSM to constrain the flux of CNO neutrinos, the pep neutrino signal was established with a significance exceeding for the first time, and the 8B flux was measured with a threshold as low as 3 MeV.
After completing the investigation of the pp-chain neutrinos, Borexino reported the first detection of solar neutrinos produced in the CNO-cycle [2020Natur.587..577B], demonstrating directly that this mode of hydrogen burning operates in stars. Profiting from a larger exposure and a better understanding of the radioactive backgrounds enabled by the unprecedented thermal stability of the detector, Borexino further improved its measurement of the CNO neutrino flux [2022PhRvL.129y2701A, 2023PhRvD.108j2005B], where the new CNO result was used in combination with existing solar neutrino results to probe solar composition.
Borexino’s physics program was not limited to solar neutrinos: its outstanding radiopurity made it a excellent detector for geoneutrinos [2020PhRvD.101a2009A] and for searches for various rare processes. Borexino data were used to constrain exotic properties of neutrinos, such as their magnetic moments [2017PhRvD..96i1103A] and non standard interactions with matter [2020JHEP...02..038B].
II.2.4 JUNO
The Jiangmen Underground Neutrino Observatory (JUNO) is a large liquid scintillator detector currently under construction in an underground laboratory with a vertical overburden of about 650 m (roughly 1800 m water equivalent) in Jiangmen city in Southern China [2015arXiv150807166A]. JUNO is located 52.5 km from two nuclear power plants, a baseline optimized for JUNO’s primary goal, the determination of the neutrino mass ordering [28, 29]. To achieve this, JUNO requires a large target mass (20 kton) and excellent energy resolution, attributes also import in solar neutrino detection. The relatively shallow overburden limits the solar neutrino program, due to cosmogenic activity, but the low threshold and large detector mass may allow measurements of day/night effects, non-standard neutrino interactions affecting the 8B spectral shape, and 7Be neutrinos, as well as providing a new solar determination of . JUNO’s reactor data will constrain to extremely high precision, allowing comparisons between the solar and reactor determinations of this parameter. To the extent that oscillation effects can be treated with greater confidence, the connections between solar neutrino flux measurements and the solar fusion reactions generating those fluxes will be sharpened [2016JPhG...43c0401A].
II.2.5 DUNE
As recommended in the 2023 P5 report [P5Report], the DUNE Collaboration seeks to construct two 10-kton fiducial liquid argon time projection chambers (LArTPC), deep underground in the Homestake mine in South Dakota [2016arXiv160102984A]. A third LArTPC will follow in Phase II, along with other upgrades. The newly developed Sanford Underground Research Facility (SURF) offers 4800-m water equivalent overburden, and will form the far site for a long baseline neutrino program utilizing a high-energy neutrino beam directed from Fermilab to SURF. The DUNE LArTPC detectors may be sensitive to measurements of the high-energy solar neutrinos via charged current interactions on argon, which offers good precision on the spectral shape and, thus, the potential to measure day/night effects [2019PhRvL.123m1803C]. The solar neutrino program at DUNE is limited primarily by backgrounds (the detector is optimized for GeV-scale physics) and energy threshold. Multiple technologies are under consideration for a fourth detector module, with the motivation of expanding the physics program, including improved sensitivity to solar neutrinos. Use of underground argon, or alternative technologies such as organic or water-based scintillators, could preserve the long baseline neutrino sensitivity while also opening up a rich program of low-energy physics.
II.2.6 Future prospects
A “hybrid” detector that could utilize both Cherenkov and scintillation light simultaneously could achieve unprecedented levels of particle and event identification and, hence, background rejection [2014arXiv1409.5864A]. The Cherenkov signature offers directional information, while the high light yield scintillation offers precision energy and vertex reconstruction. A full waveform analysis of detected light offers yet more information, based on the pulse shape of the scintillation, which is subject to species-dependent quenching effects, and the impact of the Cherenkov threshold. As a result, both the shape of the waveform and the ratio of the two signals will differ for different particle types. A full-scale detector utilizing novel scintillators along with fast and spectrally-sensitive photon detectors, such as the proposed Theia experiment [2020EPJC...80..416A], or the Jinping detector [2017ChPhC..41b3002B] could achieve percent-level precision on the CNO neutrino flux, as well as improving precision across the suite of solar neutrino measurements, such as the 8B spectral shape, and hep neutrino flux.
As experiments grow larger, and capabilities increase, there are also opportunities to apply detectors designed for other purposes to the detection of solar neutrinos. Noble liquid detectors built primarily for dark matter searches and long-baseline neutrino experiments have now reached the size and background levels to be sensitive to solar neutrinos [2024arXiv240802877A]. The detection of neutrinos from the high-energy tail of the 8B spectrum via coherent elastic neutrino-nucleus scattering (“CENS”) is particularly interesting because it is a purely neutral-current process sensitive to the total flux of active neutrinos (above the experimental threshold, about 8 MeV) from the Sun. Only the SNO experiment has had such a capability. Initial results reported in [2024arXiv240802877A] are in agreement, with limited statistical precision. Solid state detectors built purposely for observing CENS have observed neutrinos from stopped pion decay at the theoretically expected rate [doi:10.1126/science.aao0990].
The 1HH reaction ()
The cross section for the initial reaction in the pp chain (Figure 1), , is too small to be measured in the laboratory. It must be calculated from the standard theory of weak interactions.
III.1 Introduction and terminology
Near the Gamow peak energy 6 keV for temperatures characteristic of the solar core, the first and second derivatives of the astrophysical factor at zero energy, and , generate 7% and 0.5% corrections, respectively, in Taylor’s series expansion of around . Higher derivative terms are neglected in this review since they only contribute at the level. The recommended values for and will be discussed in Section III.5. Here we focus on .
At zero relative energy, can be written as [1968ApJ...152L..17B, 1969ApJ...155..501B],
where is the fine-structure constant, MeV is the proton mass, fm-1 is the deuteron binding wave number, being the proton-neutron reduced mass and the deuteron binding energy, and are the Fermi vector and axial-vector weak coupling constants. Finally, is the phase-space factor for the pp reaction with radiative corrections, is the value for superallowed transitions, and is proportional to the transition matrix element connecting the pp and deuteron states.
III.2 Adopted parameters for this review
For the phase-space factor , we use the same value as in 2011RvMP...83..195A. It comes from the value without radiative corrections, [1969ApJ...155..501B], increased by 1.62% to take into account radiative corrections to the cross section [2003PhRvC..67c5502K]. The main source of uncertainty in arises from neglected diagrams in which the lepton exchanges a weak boson and a photon with different nucleons. These diagrams are estimated to modify by , based on scaling the similar nucleus-dependent correction in superallowed decay [2003PhRvC..67c5502K]. Direct computations of these diagrams were recommended in Solar Fusion II. Here we urge again that this computation be carried out.
For , we use the value from the Particle Data Group (PDG) compilation [2022PTEP.2022h3C01W], , whose central value is larger than the 2008 PDG value, , used in 2011RvMP...83..195A by (or 1.9 ). Naively, this would lead to a increase in the central value of according to Eq. (5). Its effect will be discussed below. For , we take s from the most updated comprehensive analysis of experimental rates with the radiative and Coulomb effects corrected [2020PhRvC.102d5501H]. This value is consistent with s [2009PhRvC..79e5502H] used in 2011RvMP...83..195A with a larger error.
The dominant uncertainty in comes from the normalized Gamow-Teller (GT) matrix element . Reducing this uncertainty has been the main focus of theoretical work since 1998RvMP...70.1265A. In 1998RvMP...70.1265A, was written as , i.e. the sum of the one- and two-body current matrix elements, and , respectively, with their uncertainties estimated independently. In 2011RvMP...83..195A, two major steps had contributed to reducing the uncertainty on . The first was a much deeper understanding of the correlation between the uncertainties in and : the overall uncertainty in could be described by a universal parameter that could be fixed by a single measurement. The study of 1998PhRvC..58.1263S demonstrated this phenomenologically in the context of potential-model approaches, while later analysis via effective field theory (EFT) provided a more formal justification [2001PhRvC..63c5501B, 2003PhRvC..67e5206P]. The second step was the use of the precisely known tritium decay rate to fix this universal parameter, as first proposed by 1991PhRvC..44..619C. This has been done in both potential models [1998PhRvC..58.1263S] and in the hybrid EFT approach [2003PhRvC..67e5206P] as explained below.
In 2011RvMP...83..195A, was determined with three approaches. The first one was the potential model approach. In the most elaborate calculation for the pp fusion process, a comparison of the results for five representative modern potentials designed to accurately reproduce nucleon-nucleon scattering data was carried out [1998PhRvC..58.1263S]. After adjusting the unknown strength of the two-body exchange currents to reproduce , the variation in that otherwise would come from the choice of the phenomenological potential was largely removed. Predictions for five representative high-precision phenomenological potentials fell in a narrow interval . There were additional uncertainties in the three-body potentials and three-body currents in , of the order of , and a 0.5% uncertainty due to effective range parameters for nucleon-nucleon scattering. Hence, the recommended value from the potential model approach was .
The second and third approaches were both based on EFT. The second one was a hybrid EFT (EFT*), which used the current operators derived from EFT in conjunction with the initial and final state wave functions generated by a potential model [2003PhRvC..67e5206P]. For pp fusion, the relevant two-body current contained only one unknown low-energy constant (LEC) which parameterized the contact axial coupling to two nucleons [2003PhRvC..67e5206P]. A weakness of this approach was the mismatch between the operators and wave functions. However, it was argued that the mismatch only happened for short distance physics which could be absorbed by the LECs. Hence, when the ultraviolet cutoff was changed over a physically reasonable range, the residual cutoff dependence of physical observables provided a measure of the model dependence of the EFT* calculation. By combining the error from changing the cutoff in the range of MeV and the 0.4% higher order correction, obtained by multiplying the contribution of the highest calculated order with the small expansion parameter , 2003PhRvC..67e5206P provided the value , in perfect agreement with the one obtained within the phenomenological approach.
The third study was performed with the pionless EFT (EFT) approach. It is a framework applicable to processes with the characteristic momentum much smaller than the pion mass , such that the pion field can be “integrated out” and becomes non-dynamical (1996APS..DNP..BA03K; 1999PhRvL..82..463B; 1999NuPhA.653..386C). In this approach, all nucleon-nucleon interactions and two-body currents are described by point-like contact interactions with a systematic expansion in powers of . For all the deuteron weak breakup processes (e.g. and scattering) and their inverse processes, including the pp fusion, only one two-body current (with coupling ) is needed up to next-to-next-to-leading order (N2LO) [2001PhRvC..63c5501B]. Therefore a single measurement will fix and the rates of all such processes. This feature is shared by the other approaches discussed above. The computation of in EFT was carried out to the second order in the expansion by 2001PhRvC..64d4002K and then to the fifth order by 2001PhLB..520...87B. Constraints on from two-nucleon systems [2002PhLB..549...26B, 2003PhRvC..67b5801C] yielded .
Based on the consistent results of the above three approaches, 2011RvMP...83..195A recommended . The 1998RvMP...70.1265A value from 1969ApJ...155..501B was not re-computed in 2011RvMP...83..195A. However, with reaching a accuracy, new calculations of and , together with the full pionful chiral EFT (EFT) computations to remove the unknown systematics of the hybrid EFT, were called for in 2011RvMP...83..195A. These two challenges have been met in SF III and will be described below.
III.3 Experimental progress on muon capture of the deuteron
In 2011RvMP...83..195A it was also recommended to carry out the experimental determination of the muon capture rate on deuteron, as proposed in the MuSun experiment [2003nucl.ex...4019K, 2017APS..DNP.CG007S]. This quantity could be used to constrain without the need to rely on the three-body calculation of . The theoretical calculations for the muon capture rate on deuteron have been carried out in EFT [2005PhRvC..72f1001C] (it is possible to impose neutron energy cut to isolate the low-energy neutron events so EFT is applicable [2003nucl.ex...4019K]), chiral hybrid EFT* [2002PhLB..533...25A], phenomenological potential model [2011PhRvC..83a4002M], and recently in EFT (2012PhRvL.108e2502M, 2018PhRvL.121d9901M; 2018PhRvC..98f5506A; 2023FrP....1049919C; 2023PhRvC.107f5502B; PhysRevC.109.035502). However, the MuSun result is yet to be released. On the other hand, we should notice that the muon capture processes happen at a rather large momentum transfer compared to pp fusion. The momentum transfer dependence of the single nucleon axial coupling constant , with , has been recently studied by 2018RPPh...81i6301H, who have provided an experimental determination for the axial charge radius, given by fm2, where is defined by the relation for small . The % uncertainty on has an impact on the ability of the MuSun experiment alone to directly constrain [2018PhRvC..98f5506A, 2023FrP....1049919C, 2023PhRvC.107f5502B, PhysRevC.109.035502].
In the next few years, lattice Quantum Chromodynamics (QCD) calculations of are expected to reduce the uncertainty by a factor of two or more (see 2023slft.confE.240M for the most recent review of the lattice results). Currently, lattice QCD results are consistent with 2018RPPh...81i6301H on , often with comparable uncertainties. At larger ( GeV2), there is a growing tension between lattice QCD predictions of and the phenomenological determination from older neutrino-deuterium bubble chamber data [2016PhRvD..93k3015M] with the lattice QCD results yielding a 30% larger neutrino-nucleon cross section over a large range of [2022ARNPS..72..205M].
In any case, it is evident that accurate experimental determinations of muon capture rates on the deuteron and other light nuclei, which can be addressed theoretically ab initio approaches, represent fundamental tests for the theoretical approaches themselves, either within EFT or EFT, and might be able to provide the necessary experimental information to fix the unknown parameters of the theory.
III.4 Progress in calculations since SF II
Below we summarize the calculations performed after 2011RvMP...83..195A using different approaches.
III.4.1 EFT
The pioneering work of 2013PhRvL.110s2503M used the next-to-next-to-next-to-leading-order (N3LO) chiral two-nucleon potential [2003PhRvC..68d1001E, 2011PhR...503....1M] augmented with higher order () two-photon and vacuum-polarization electromagnetic interactions. These corrections reduced by % mainly due to the vacuum-polarization-induced pp wave function distortion. This was consistent with the 0.84% first found in the potential model calculation of 1998PhRvC..58.1263S. This correction was also included in the EFT* calculation of [2003PhRvC..67e5206P]. The relevant LECs were fitted to reproduce the binding energies, magnetic moments, and GT matrix element in to obtain , with the -wave initial state contributing at , which was about the accuracy level of the calculation [2013PhRvL.110s2503M]. However, using EFT at next-to-leading-order (NLO), 2019PhRvC.100b1001A later found that -wave only contributed at the order of . Re-examining the computer programs, 2013PhRvL.110s2503M found an error in the determination of one of the -wave reduced matrix elements (associated with the longitudinal multipole operator). Consequently, 2019PhRvL.123a9901M reported in the Erratum that .
In more recent work, 2016PhLB..760..584A used chiral interactions and consistent currents up to N3LO (called next-to-next-to-leading-order (N2LO) in the original literature since the second order vanished) and developed a robust procedure for the error quantification. In particular, they analyzed a family of 42 interactions [2016PhRvX...6a1019C] with 7 different cutoff values from 450 to 600 MeV. The 26 LECs were fitted to 6 different pools of input data including and scatterings, as well as the binding energies and charge radii of 3H and 3He, the quadrupole moment of 2H, and . This thorough study yielded .
In both of the EFT calculations mentioned above, a widely used relation first proposed by 2009PhRvL.103j2502G that linked the two-body axial current LEC with the LEC from the vertex was employed. However, later 2018PhRvL.121d9901M found that there was a factor missing in this - relation which was then acknowledged in the Erratum of 2019PhRvL.122b9901G. Fortunately, this error was unimportant in muon capture on deuteron [2018PhRvL.121d9901M] and it only affected at the level [2023arXiv230403327A].
In the most recent and comprehensive EFT study, 2023arXiv230403327A compared the above calculations of 2019PhRvL.123a9901M and 2016PhLB..760..584A in detail. In addition to the increase of from using the correct - relation, updating the input parameters to their most recent values increased by in both calculations, mainly due to the increase of from its 2011RvMP...83..195A value mentioned above. Furthermore, 2019PhRvL.123a9901M received a increase from removing the truncation error of the basis functions which effectively cut off the long distance part of the wave functions, as 2017PhRvC..95c1301A advocated. After these corrections, the difference between 2019PhRvL.123a9901M and 2016PhLB..760..584A was reconciled and the combined result was found to be [2023arXiv230403327A]
2023arXiv230403327A obtained consistent values of using four different EFT models for the nuclear interaction. This value is also in agreement with the result obtained in EFT by 2022arXiv220710176D (see below). In estimating the order-by-order convergence, however, 2023arXiv230403327A used one of the EFT models, which is able to nicely reproduce the deuteron properties already at leading order. Therefore, the EFT error of Eq. (6) is likely to be an underestimate, warranting the enlarged error advocated in Section III.4.4.
Finally, we would like to mention the work of 2022PhRvC.106e5501L, where the power counting of the EFT weak current operator involved in the pp reaction is revisited using renormalization group (RG) invariance as the guideline. In particular, it is argued that the contact two-body axial current proportional to the LEC must appear one order lower than assessed by naive dimensional analysis. Then it can be shown that RG invariance is fulfilled at N2LO. However, the estimate for obtained by 2022PhRvC.106e5501L does not use a value for obtained by fitting , but extracted in order to match the value of obtained in 2011RvMP...83..195A. Therefore, we will not consider the work of 2022PhRvC.106e5501L in the present evaluation.
III.4.2 EFT
The universal two-body current coupling was determined using in EFT for the first time by 2019PhRvC.100e5502D. This calculation was carried out up to NLO using the dibaryon formulation of 2001NuPhA.694..511B, which partially resumed higher order effective range contributions to improve the convergence. Then, this result was used by 2022arXiv220710176D, with updated input parameters, to obtain
where the electromagnetic correction was not calculated but assumed to be the same as the potential model value, [1998PhRvC..58.1263S]. While this number was not model independent, the model dependence was believed to be well below the assigned error.
The small error assigned to this NLO result has been justified by drawing an analogy from the corresponding electromagnetic processes. Using the same approach, the matrix element at threshold was predicted at NLO within 0.5% to the experimental value after the electromagnetic two-body current was fit to the magnetic moments of 3He and 3H. This indicates that the contribution of the three-body current at N2LO is small in this electromagnetic case. The weak sector is shown to follow the same operator structure and hence provides support for the calculation procedure and the uncertainty estimate.
It is important to note that the reason that the three-body current is an N2LO effect in EFT is related to the non-trivial renormalization of the non-derivative three-body contact interaction, which shows up at LO to absorb the cutoff dependence of Feynman diagrams. The subleading two-derivative three-body contact interaction is expected to show up at N2LO. This interaction, combined with the one-body current, renormalizes the three-body current. Hence, the three-body current should also appear at the same order, N2LO. However, if the non-derivative three-body contact interaction were counted as higher order, such as N3LO, as in EFT (because cutoff independence is not strictly enforced order by order in EFT), then the three-body current would contribute at much higher order. Although the EFT power counting indeed yields good convergence in the expansions, it is unsatisfactory that one can not remove the cutoff dependence at each order of the expansion. In addition, the uncertainty estimate of the EFT still lacks a broad inspection of the specific nuclear EFT potential implementation. As a consequence, we consider the EFT 1.5% error a better estimate for the theoretical uncertainty.
III.4.3 Lattice QCD and Lattice EFT
Ideally lattice QCD would provide a first principles prediction of the pp fusion rate and the GT matrix element of , for both pure QCD and with QED effects incorporated. However, such calculations are quite challenging and not yet available at the required precision. A proof-of-principle calculation was carried out by 2017PhRvL.119f2002S using a background field method to determine both the pp fusion GT matrix element and . This exploratory calculation utilized a single pion mass at the SU(3)-flavor–symmetric point with MeV, a single and relatively coarse lattice spacing of fm, and a single volume. The calculation was performed under the assumption that the two- and three-nucleon systems were deeply bound. Without this assumption, matrix elements computed in the finite volume can be significantly different from those in infinite volume, due to Lellouch-Lüscher factors [2001CMaPh.219...31L] that lead to power-law finite volume corrections. These can range from the few-percent level to [2013PhRvD..88i4507B, 2015PhRvL.115x2001B].
More recent lattice QCD calculations have found that two-nucleon systems at heavy pion masses are in fact not bound (2016JHEP...10..101I; 2019PhRvD..99g4505F; 2021PhRvC.103a4003H; 2021arXiv210810835A). These efforts, which employ interpolating operators more sophisticated than those of 2017PhRvL.119f2002S, suggest there could be large systematic uncertainties affecting the conclusions of 2017PhRvL.119f2002S stemming from misidentification of the spectrum and inaccurate Lellouch-Lüscher factors. In addition, 2021PhRvL.127x2003G found that the two-baryon spectrum may be particularly sensitive to discretization effects, which would also have important implications for the continuum extrapolation of the matrix elements. These issues are discussed in some detail in 2022FBS....63...67T. Finally, results with pion masses MeV are needed for accurate extrapolation to the physical pion mass.
Given these unresolved systematic issues, the result of 2017PhRvL.119f2002S is not included in the present evaluation – even though the extracted value is consistent with our recommended range quoted in Section III.4.4.
The lattice EFT computation of 2015PhLB..741..301R performed the pp fusion calculation by implementing EFT on a spacetime lattice. The purpose of this leading-order study was to demonstrate that lattice EFT could reproduce the infinite volume and continuum result of EFT such that it could be applied to various reactions of astrophysical interest in the future. Therefore, for , this result is considered as a subset of the EFT calculation.
III.4.4 Final Recommendation of
The above discussions show that determinations of from EFT in Eq.(6), EFT in Eq.(7), and lattice QCD (although with unquantified systematics) are all consistent with each other. Furthermore, these values are also consistent with the recommended value of 2011RvMP...83..195A, provided the central value is increased by 555To know the precise shift requires an explicit calculation. However, if the shift is within the range of 0.8-1.0 , the averaged value and remain the same within the significant digits. to to account for the update. Averaging this value with the EFT value in Eq.(6) and the EFT value in Eq.(7) yields with per degree of freedom to be . This shows that the EFT, EFT and 2011RvMP...83..195A estimates of are all mutually consistent.
In addition, we would like to advocate adding an additional correlated error to account for any input that would tend to move all results in a coordinated way. For example, from 2011RvMP...83..195A to SF III, we experienced the shift due to the update of the PDG value of . It is not inconceivable that or other input parameters or physics could change again by similar amounts: the large PDG inflation factor of 2.7 reflects the tension that continues to exist among measurements [2022PTEP.2022h3C01W]. Therefore, assigning an additional correlated error seems reasonable. This is also in line with the subtleties discussed in Secs.III.4.1 and III.4.2 which call for an enlarged error. Therefore, our final recommended value for is
where we have added the correlated and uncorrelated errors linearly to be conservative.
III.5 Progress in and
Using EFT, 2013PhLB..720..385C computed analytically with all partial waves included up to N2LO. The Fermi matrix element only contributed at the level and was neglected compared with the GT matrix element. The energy dependence of the phase factor of Eq. (5) was the dominant effect in and , the energy dependence of pp scattering was subdominant, while the contribution was much less important in these derivatives than in . Therefore, these derivatives could be predicted more reliably than . Furthermore, the derivatives were computed analytically and were free from errors of fitting to a polynomial. The result was MeV-1 and MeV-2.
In EFT, for keV was fit to polynomials of [2019PhRvL.123a9901M]. Depending on using a quadratic or quartic fit, changed from 12.23 to 10.82 MeV-1 and changed from 178.4 to 317.4 MeV-2. 2016PhLB..760..584A used cubit fit and keV to obtain =10.84(2) MeV-1 and =317.8(13) MeV-2. Recently 2023arXiv230403327A repeated the calculation of 2019PhRvL.123a9901M with the same energy range and cubic fit as 2016PhLB..760..584A, obtaining a consistent result with 2016PhLB..760..584A: MeV-1 and MeV-2. We will take these as the recommended values from EFT.
Although the face values of and from EFT and EFT look quite different, we would like to remark that they actually agree on better than 0.1% below the keV Gamow peak. For massive stars with central temperatures keV, the agreement is better than 0.8% to second order in the derivatives and 0.5% if MeV-3 is included in the EFT result. Hence, we take the average of the EFT and EFT results as the recommended value:
Finally, we comment on the work of 2019PhRvC.100c5805G which performed a study of in a wide energy range. The main focus of this work was to perform a proper energy-dependence analysis of the pp process, in order to reliably extract and . However, the calculation was performed within a phenomenological approach, using a quite simplified model for the nuclear currents (i.e. not including two-body currents, which are well known to be significant) and structure (i.e. neglecting the -wave components in the deuteron wave function). They found results for and compatible with those of Eq. (9). However, since these values were not obtained with state of the art calculations, they have not been considered in the determination of and .
The reaction ()
IV.1 Introduction
The 2H(,)3He reaction is the second step in the pp chain (Figure 1). Compared with the reactions mediated by the weak interaction, this reaction occurs much more rapidly. Consequently, on the time scales relevant to solar energy generation, deuterium is effectively converted to 3He instantaneously and thus it is only sensitive to the value of the reaction and not the uncertainties of the rate.
However, the 2H(,)3He reaction plays an important role in the development of protostars, because the onset of deuterium burning slows down the protostars contraction and heating, increasing their lifespan. Accurate knowledge of the 2H(,)3He reaction rate, particularly within the few keV range corresponding to the Gamow peak in protostars, is vital for modeling protostellar evolution effectively [1988ApJ...332..804S].
Another astrophysical scenario where the 2H(,)3He reaction plays a key role is Big Bang Nucleosynthesis (BBN), responsible for the production of light elements during the first few minutes of the Universe. Among these elements, deuterium is an excellent indicator of cosmological parameters because its primordial abundance is the most sensitive to the baryon density and critically depends on the radiation density of the early Universe, see for example a review of BBN [2016RvMP...88a5004C]. The reactions involved in the synthesis of deuterium are: production via the well known )2H process and destruction via the 2H(2H,)3He, 2H(2H,)3H and 2H(,)3He reactions [2020JCAP...03..010F].
Since the comprehensive review performed in 2011RvMP...83..195A, there have been both new measurements as well as advances in the theoretical and phenomenological analysis of the 2H(,)3He reaction. The new experimental results have been determined with accelerator-based measurements of the cross section [2019EPJA...55..137T, 2020Natur.587..210M, 2021PhRvC.103d5805T], as well as plasma-based, intertial confinement fusion measurements [2020PhRvC.101d2802Z, 2022FrP....10.4339M] and are reviewed in Section IV.2 below. On the theoretical side, there have been advances in ab initio calculations, where nucleons are the fundamental degrees of freedom interacting among themselves and with the external electromagnetic probe, see for example [2023FrP....1129094E]. Finally, Bayesian analysis methods have been used to model the energy dependence of the S-factor, starting from the ab initio predictions and applying a polynomial approximation to it.
We review these updates and provide recommended values and uncertainties of the S-factor over the energy range of interest for Solar fusion, based upon a Bayesian averaging of various models.
IV.2 Data sets used for the present review
The 2H(,)3He reaction has a Q value of 5.5 MeV and proceeds through the radiative capture of a proton on deuterium. Different experimental approaches were followed to measure its cross section. 2019EPJA...55..137T irradiated deuterated titanium targets with a proton beam and detected the -rays with two high-purity germanium (HPGe) detectors placed at different angles. The final S-factor is provided at four energies in the 47 - 210 keV range, with approximately 15% statistical uncertainty. More recently, the LUNA Collaboration performed a measurement in the energy range of keV at the underground in the Gran Sasso Laboratories, exploiting the six orders of magnitude suppression of the cosmic radiation background [2021FrASS...7..119Z, 2018IJMPA..3343010C]. A windowless deuterium gas target was used and the -rays emitted by the 2H(,)3He reaction were detected by a large HPGe detector at 90∘ with respect to the beam axis. Great care was taken to minimize all sources of systematic uncertainties in the S-factor overall uncertainty at the 3% level [2020EPJA...56..144M, 2020Natur.587..210M]. These new results provided stringent constraints on cosmological parameters obtained by comparing the precise primordial deuterium abundance predictions of the standard BBN model with astronomical observations [2018ApJ...855..102C]. A deeper discussion of the LUNA results and their implications can be found in [2021MNRAS.502.2474P, 2021JCAP...04..020P, 2021JCAP...03..046Y, 2021ApJ...923...49M] 666After the closing of the present review, the 2020Natur.587..210M data were re-analyzed to extract the experimental 2H()3He -ray angular distribution [2024PhRvC.110c2801S]. The data confirm the ab initio angular distributions used in the original analysis by 2020Natur.587..210M.
Finally, a new measurement was performed at the Helmholtz-Zentrum Dresden-Rossendorf in the 300-1000 keV energy range using implanted deuterium targets on tantalum backings and two HPGe detectors [2021PhRvC.103d5805T]. The resulting S-factors show tension with the analysis of the LUNA results extrapolated to keV [2020Natur.587..210M]. However, they are affected by large systematic uncertainties. A new measurement of the reaction is planned at the Felsenkeller laboratory in Germany [2019sone.conf..249B].
In addition to the new accelerator-based results, two recent sets of measurements [2020PhRvC.101d2802Z, 2022FrP....10.4339M] have also been performed using the inertial confinement fusion plasma-based platform [2017PhPl...24d1407G, 2022FrP....10.4339M, 2023FrP....1180821G], which has recently begun to be exploited for this type of work [e.g., 2016PhRvL.117c5002Z, 2017NatPh..13.1227C]. Both measurements, which were performed at the OMEGA laser facility (see section LABEL:subsec:WG8:Plasma), used laser-driven implosions of spherical plastic-shell capsules filled with H2D2 gas, and measured the emitted -rays using a gas-Cherenkov-detector [2014RScI...85kE124H] that was calibrated applying the technique described in 2019RScI...90l3504Z. The initial experiment [2020PhRvC.101d2802Z] obtained good statistics at an energy of 16 keV by making several repeated measurements, with a final statistical uncertainty of 6% and a systematic one of 17% (dominated by uncertainty in the absolute calibration of the detector). The second experiment [2022FrP....10.4339M] obtained data at three different energies in the 17-37 keV region, with comparable systematic uncertainty, but with larger statistical uncertainty due to fewer repeated measurements. The results obtained on this unique platform agree within error bars with the accelerator-based measurements.
IV.3 Theoretical studies
Nuclear reactions of astrophysical interest in general, and the 2H(,)3He in particular, are of great importance in nuclear theory because the available experimental data can be used to test the adopted theoretical framework. The 2H(,)3He reaction has the great advantage of involving only nuclear systems, and can be addressed with a microscopic ab initio study. This means that the nuclear systems involved in the process are viewed as made up of nucleons, interacting among themselves and with the external electromagnetic probes. Within such an approach, the following ingredients are essential for the calculation: (i) realistic models for the nuclear interactions and currents, possibly rooted in Quantum Chromodynamics (QCD); (ii) a numerical technique able to solve the -body bound and scattering state problem, including the Coulomb interaction without approximation. Such a technique is usually referred to as an ab initio method. The agreement (or disagreement) between ab initio theoretical predictions and experimental data represents a validation (or indicates the necessity of improvement) mostly for ingredient (i), i.e. are the models of the nuclear interactions and currents accurate within the precision specified by the ab initio method? This is why few-nucleon reactions can be used as an “ideal” laboratory, where the ab initio framework can be stringently tested in systems under more theoretical control.
The most recent ab initio calculation of the 2H(,)3He reaction is that of 2016PhRvL.116j2501M. Here the pair-correlated Hyperspherical Harmonics ab initio method was used to calculate the initial scattering and final bound state wave functions (see 2008JPhG...35f3101K and 2020FrP.....8...69M for details). The nuclear interaction model adopted to describe the nuclear state in 2016PhRvL.116j2501M consists of a two-nucleon term, the Argonne (AV18) potential [1995PhRvC..51...38W], augmented by a three-nucleon contribution, the Urbana IX (UIX) potential [1995PhRvL..74.4396P]. The AV18 potential can reproduce the large two-nucleon database with a /datum [1995PhRvC..51...38W], while the combination AV18/UIX can describe quite accurately the properties of 3He, the spectra of light p-shell nuclei [2001ARNPS..51...53P], and scattering observables (see for instance 2002PhRvC..65c4002W). The electromagnetic current operator used in 2016PhRvL.116j2501M includes, in addition to the non-relativistic one-body operator, two- and three-body terms required by gauge invariance in a system of interacting particles. These terms were constructed in 2005PhRvC..72a4001M. The model was then tested against various electromagnetic observables, to access the quality of its predictions. As a potential model, however, there is no systematic procedure for assigning uncertainties for observables whose values are unknown.
The results of 2016PhRvL.116j2501M have been found to be about 10% higher than the experimental data of 2020Natur.587..210M. Given the lack of a procedure for quantifying errors in calculations based on phenomenological interactions and currents, it is difficult to access the significance of this discrepancy. This leads us to make several recommendations to the theory community. First, the 2H(,)3He reaction should be studied within the framework of chiral effective field theory, which has reached a degree of accuracy and predictive power comparable to potential-based phenomenology. This approach is formally ab initio and rooted in QCD, and as an operator expansion will provide an estimate of the theoretical uncertainty. Work along this line is currently underway. Second, the angular distribution of the 2H(,)3He capture reaction, and possibly also polarization observables, should be both calculated and measured. This would be valuable even if measurements were limited to higher energies, where they are less difficult. Such measurements would provide a further test of the predictive power of the theory.
IV.4 Phenomenological and Bayesian Analyses
In the energy range of astrophysical interest, the 2H(,)3He reaction does not have any resonance or coupled channels that can give rise to non-trivial energy dependence, and the S-factor can be modelled by a low-order polynomial in energy [1967ARA&A...5..525F]. Moreover, the recent results from LUNA [2020Natur.587..210M], combined with previous measurements, place stringent constraints on the cross section, reducing the uncertainty over the range of interest to Solar fusion and BBN [2002NuPhA.706..203C, 2020Natur.587..210M, 2021MNRAS.502.2474P, 2021JCAP...04..020P, 2021JCAP...03..046Y].
A significant change since 2011RvMP...83..195A has been the wide adoption of Bayesian analysis methods to evaluate thermonuclear reaction rates and provide more rigorous uncertainty estimates in a statistical sense [2016ApJ...831..107I]. This approach has been followed in the analysis presented here, which closely follow work on the 2H(,)3He reaction reported in 2021ApJ...923...49M, and references therein. The mathematical details of the Bayesian method will be found in Appendix LABEL:AppendixWG4: Here we focus on the application to 2H(,)3He by the authors cited above.
Both the ab initio prediction from 2005PhRvC..72a4001M as well as a third-order polynomial in the energy were used to constrain the data, where we define an -order polynomial as
As noted in the literature, the energy dependence of 2005PhRvC..72a4001M is more reliable than the absolute normalization. Furthermore, the updated prediction in 2016PhRvL.116j2501M, as noted above, is larger than that of 2005PhRvC..72a4001M. Therefore, 2021ApJ...923...49M modeled the prediction of 2005PhRvC..72a4001M, denoted , as
where and are an unknown scale factor and offset to be determined in the analysis. While was not determined with theoretical uncertainty, model uncertainty is introduced through the parameters and . The resulting mean values and uncertainties of determined from Eq. (11) were found to be comparable to those with the third-order polynomial.
IV.5 Summary and Recommendations
| Model | logGBF | weight |
| ab initio, | 1478.7 | 0.585 |
| ab initio, | 1478.3 | 0.414 |
| 3rd order polynomial | 1470.2 | 1.2 |
| 4th order polynomial | 1469.2 | 4.6 |
| 5th order polynomial | 1468.2 | 1.6 |
| 6th order polynomial | 1464.1 | 2.9 |
In this review, we take the further step of applying Bayes Model Averaging to a number of reasonable models that describe the data, which systematically captures additional theoretical “model selection uncertainty”. Detailes of the Bayesian analysis and model averaging procdure are presented in Appendix LABEL:AppendixWG4. The models we explore include Eq. (11) with both and , as done in 2021ApJ...923...49M, as well as setting , and then we consider polynomials in energy of order .777The implementation of the Bayesian analysis presented here is different than that of 2021ApJ...923...49M but it was verified to produce the same results when the same implementation of the systematic uncertainties was made.
With such a Bayes model averaging, we can quantitatively compare and contrast the polynomial parameterizations of 2H(,)3He along with ab initio results predicted by 2005PhRvC..72a4001M, 2016PhRvL.116j2501M, including the extra variance that arises from this set of reasonable models, e.g., using polynomials of different order, as well as the phenomenological model of Eq. (11). Results are listed in Table 3. The first column lists the model (i.e., polynomial or ab initio), the second column the natural logarithm of the Gaussian approximation to the BF, and the third column the corresponding weight in the model averaging. From this model averaging, the resulting prediction for the 2H(,)3He S-factor at a few representative energies is provided in Table 4, where the first uncertainty arises from the first term in Eq. (LABEL:eq:BMA_Var) and the second uncertainty is from the second term, which we denote as model selection uncertainty. The resulting Bayes model average prediction of the S-factor over the entire kinematic range considered is depicted in Fig. 5, with the gray band representing the 68% coverage probability. For this particular reaction, it is interesting to note that the Bayesian analysis strongly favors the phenomenological models of Eq. (11) over the polynomial approximations. One reason for this might be that the scaled ab initio models have only one or two free parameters, and the energy dependence of given in 2005PhRvC..72a4001M, 2016PhRvL.116j2501M is sufficient to accurately describe the various data sets. On the other hand, polynomial approximations are disfavored as they require more parameters to capture the energy dependence. Also of note, in the model of Eq. (11) with and using 2016PhRvL.116j2501M for , the scale factor is given by
indicating that the prediction in 2016PhRvL.116j2501M overestimates the 2H(,)3He data by 7.9%, consistent with expectations noted above. In comparison, the third order polynomial fit predicts values of that are higher at and lower at keV as compared to those in Table 4.
The analysis in Table 3 and Fig. 5 can be reproduced with the code linked in Table 4. The code also can provide a prediction of the mean value and uncertainty of at any energy over the same kinematic range.
| keV | 10 | 20 | 40 | 80 | 91 | 100 | 120 |
| 2.028(51)(9) | 2.644(60)(8) | 3.276(70)(7) | 4.579(94)(5) | 7.31(15)(0) | 8.11(16)(0) | 8.77(18)(0) | 10.24(21)(0) |
The reaction )
The 3He(3He,)4He reaction terminates the pp-I chain. The ratio of its rate to that of the 3He()7Be reaction controls the branching to the pp-II and pp-III chains, so that in 1998RvMP...70.1265A increasing was discussed as a potential solution to the Solar Neutrino Problem. Subsequent experiments, notably a very low energy measurement at the LUNA 50 kV accelerator deep underground in Gran Sasso [1999PhRvL..82.5205B] and a complementary experiment at somewhat higher energies [2004PhRvC..69a5802K] ruled out such an increase in , as summarized in 2011RvMP...83..195A.
V.1 Shape of the particle spectrum
Since no new absolute measurements of have been reported since 2011RvMP...83..195A, we consider the same four experiments [1999PhRvL..82.5205B, 1998PhRvC..57.2700J, 2004PhRvC..69a5802K, 1987NuPhA.467..273K]. However, there is new information on the energy spectrum of the emitted protons. This spectrum has recently been measured using Inertial Confinement Fusion (ICF) plasmas for a Gamow peak energy of 165 keV [2017PhRvL.119v2701Z]. The results show significant structure, indicating the presence of a sequential reaction mechanism passing through the unbound ground state of . This spectrum is important for all of the measurements since they determined cross sections by detecting only the protons above an energy threshold.
The efficiency correction by which 1987NuPhA.467..273K account for the threshold is not well documented, but the more recent measurements [1999PhRvL..82.5205B, 1998PhRvC..57.2700J, 2004PhRvC..69a5802K] utilized the genbod event generator [James68-GENBOD]. It employs a simple reaction model without final state interactions, the Pauli principle, or Coulomb effects that are important near the spectrum endpoint. The only angular correlations are those required by energy and momentum conservation, and together these simplifications give simple ellipses for the singles energy distributions of the emitted nuclei. Published proton spectra obtained using accelerator beams do exist (Fig. 2 of 1974PhRvC...9..805D and Fig. 3 of 1987NuPhA.467..273K), and although they are not corrected for instrumental effects, they do not appear to be well described by ellipses.
The 3He(3He,)4He reaction may proceed via several sequential mechanisms, including (with in its ground or first excited state) and (where the di-proton is two correlated protons in a singlet state) [2015PhRvC..92a4003B]. More complicated three-body decay channels, sometimes called direct decays, are also possible. Only a relatively narrow intermediate state (here, the 1 MeV wide 5Li ground state) could produce a peak in the energy spectrum, and precise classification of the reaction mechanism is in general both experimentally and theoretically ambiguous. When coincident detection of reaction products is used [as in 1999PhRvL..82.5205B], possible angular correlations between the reaction products also matter. A reaction through the ground state of would emit the second proton preferentially either along or opposite the direction of the first proton, while di-proton emission would tend to send both protons in the same direction [2015PhRvC..92a4003B].
The solid curves in Fig. 6 show two different R-matrix models of the proton energy spectrum at keV, fitted to the spectrum measured by 2017PhRvL.119v2701Z for MeV. The measured spectrum differs significantly from the elliptical spectrum of genbod. Based on their Fig. 6, the experiment of 1998PhRvC..57.2700J had a detection threshold of about 5 MeV proton energy. The measurements by 1999PhRvL..82.5205B required a coincidence between two detectors, with a detection threshold of 2 MeV proton energy in each detector. Finally, Fig. 19 of 2004PhRvC..69a5802K indicates a detection threshold of about 4 MeV proton energy. We have estimated their sensitivities to the assumed proton spectrum by integrating the curves shown in Fig. 6 above energy thresholds of 2, 4, and 6 MeV. These integrals vary from 3% below to 6% above the result from an elliptical spectrum, depending on the specific threshold and the assumed spectrum.
V.2 Recommendation
Based on these investigations, our recommended fit is the one from 2011RvMP...83..195A, but with an additional 4% systematic uncertainty typical of the 3% to 6% corrections to the extrapolated proton spectra estimated above to take the uncertain spectral shape into account. This gives
Only the larger constant term in the uncertainty differs from the previous evaluation, and it is due to the spectral shape discussed above.
New measurements of this reaction could provide more accurate absolute cross sections by using lower energy thresholds and improve our understanding of the proton energy spectra and angular correlations. The analysis of new measurements should use Monte Carlo simulations considering a variety of plausible energy spectra and angular correlations, in order to estimate the sensitivity to these effects. We note finally that the proton energy spectrum may depend on , as this has been found to be the case for neutrons from the mirror reaction [2018PhRvL.121d2501G].
The reaction ()
The 3He()7Be reaction proceeds via non-resonant capture to the ground and 429 keV first excited states of 7Be. It has been studied experimentally using three main methods, with references given in subsections VI.1 and VI.3: First, by detecting at least two of the three -rays from the reaction and taking the angular correlation with the alpha beam direction into account (prompt- method). Second, by detecting the induced 7Be [2002NuPhA.708....3T] radioactivity (activation method). Third, by counting the 7Be recoils (recoil method).
VI.1 Previous recommendation in SF II
The 3He()7Be S-factor recommended in 2011RvMP...83..195A was developed in multiple steps. First, a model was selected for the shape of , based on existing nucleon-level calculations [1986NuPhA.460..559K, 2001PhRvC..63e4002N]. A rescaling of the curve was fitted to data, but only at 1.002 MeV center-of-mass to minimize the role of short-distance physics in the models.
In 1998RvMP...70.1265A, a possible systematic discrepancy between data from the two methods previously used was discussed. By the time of 2011RvMP...83..195A, in-depth studies from two groups using both activation and prompt methods were available, from LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G] and Seattle [2007PhRvC..76e5801B] groups. These studies did not find any discrepancy between activation and prompt- data in direct comparisons. 2011RvMP...83..195A opted to limit the fitting to the data by the activation and recoil methods. The prompt- data were left out because of their somewhat larger common mode errors, concerns about how well the -ray angular distribution was known, and avoidance of the correlated errors between activation and prompt data from the same experiment.
In 2011RvMP...83..195A, only data published after 1998 were included. The 2011RvMP...83..195A recommended value was , based on the data by the Weizmann [2004PhRvL..93z2503S], LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G], Seattle [2007PhRvC..76e5801B], and ERNA [2009PhRvL.102w2502D] groups.
VI.2 Theory progress on
Significant theoretical work on this reaction has occurred since 2011RvMP...83..195A, but the basic understanding of its mechanism remains unchanged from the 1960s [1963PhRv..131.2582T]. It is dominated by external non-resonant capture into the two bound states, and most of the dipole strength at low energy arises beyond the range of nuclear interaction. The external capture part of the cross section is determined by the ANCs of the bound states and by scattering phase shifts; near threshold, most of the strength lies at –20 fm [2011PhRvL.106d2502N]. In models with explicit wave functions the shorter-range strength largely cancels out due to effects of nucleon-exchange antisymmetry. All models feature a shallow minimum of near 1.25 MeV, where capture from -waves becomes comparable to that from -waves.
The first fully ab initio calculation [2011PhRvL.106d2502N] appeared just after the 2011RvMP...83..195A analysis concluded; it used the fermionic molecular dynamics (FMD) method and a softened representation of the Argonne interaction [1995PhRvC..51...38W]. It agrees well with both the scale and the energy dependence of the modern S-factor data, and it is very close to the energy dependence assumed in 2011RvMP...83..195A. Neff’s elastic-scattering phase shifts also agree well with experiment. Notably, Neff found that the dipole strength distribution departs significantly from pure external capture at 3He-4He separations as large as 9 fm, compared with 3-4 fm in potential models.
Another ab initio model [2016PhLB..757..430D, 2019PhRvC.100b4304V] has also appeared, based on the no-core shell model with continuum (NCSMC) method and the chiral interaction of 2003PhRvC..68d1001E, “softened” by the similarity renormalization group (SRG) procedure. Freedom to choose the SRG stopping point allowed exact reproduction of the 7Be breakup energy, which corrects for the main effect on the external capture of omitting the three-nucleon potential. The results agree well with the overall scale of the modern data, but their energy dependence and phase shifts depart from experiment, possibly due to the omission of explicit three-body forces.
An important advance since 2011RvMP...83..195A has been the application of halo effective field theory (halo EFT) methods to astrophysical capture reactions [2002NuPhA.712...37B, 2018EPJA...54...89H, 2020EPJA...56..166P, 2020JPhG...47e4002Z]. EFTs are valid for systems with a natural separation between the momentum scales probed in low-energy experiment (e.g., corresponding to a binding energy or a projectile energy) and the much larger momentum scale where the low-energy degrees of freedom are no longer valid: the word “halo” here refers to nuclei with only shallow bound states. In halo EFT the only degrees of freedom are the initial- and final-state nuclei treated as point particles, plus photons. For , MeV corresponds to the momentum needed to separate a proton from 3He or to excite 4He; the thresholds to separate the two 7Be bound states into 3He and 4He correspond to momenta MeV and 60.9 MeV.
Given a sufficient separation of scales, one constructs a Lagrangian that respects the system’s symmetries and known qualitative features systematically, organized by powers of . This series can be truncated (at leading order or LO, next-to-leading order or NLO, next-to-next-to-leading-order or NNLO, etc.), and the precision of the resulting theory depends on the sizes of the omitted terms. The precision of a calculation and the energy where it breaks down can be estimated by assuming coefficients of the first omitted term to have a “natural” size. Coupling constants of the Lagrangian must be fitted to data, and consistency of the power counting scheme (identification of powers of for the main operator terms of a system) is tested by whether the fitted constants have natural sizes. Low-order coupling constants in a halo EFT can often be identified with familiar quantities like ANCs, scattering lengths, and effective ranges; the description of elastic scattering in halo EFT reproduces the (Coulomb-modified) effective-range expansion [2022JPhG...49d5102P, 2000NuPhA.665..137K, 2008NuPhA.809..171H]. When a separation of scales exists, halo EFT is well-suited to data extrapolation because it avoids the tacit, hard-to-test, and unavoidable prior assumptions present in the short-range parts of models based on explicit wave functions.
Two groups have studied in halo EFT [2018EPJA...54...89H, 2020EPJA...56..166P, 2020JPhG...47e4002Z]. The sizes of the Coulomb interaction and the large -wave scattering length make the correct power counting tricky to establish for this system. The work of 2018EPJA...54...89H and 2020EPJA...56..166P includes a careful examination of possible power-counting schemes and strict adherence to a power counting once established, up to terms of NLO or NNLO, respectively, in two different power-countings. These authors fitted EFTs to S-factor data, and they examined effects of including or excluding elastic scattering constraints in their fits. When they included scattering phase shifts from 1972NuPhA.195..241B in their fits, they concluded that the large-scattering-length power counting of the NLO theory was favored. This result promotes two-body currents (i.e., contributions not equivalent to external non-resonant capture) to leading order in the theory. The -wave contribution that becomes important above 1 MeV first appears at NLO. In addition to data-fitting errors, these theories have errors estimated to be 10% from EFT truncation.
The work of 2020JPhG...47e4002Z followed a different approach to power counting up through NLO. At MeV, their derived expression for is essentially the same as that from the 2020EPJA...56..166P NLO theory. 2020JPhG...47e4002Z also developed some ad hoc (i.e., not systematically developed) higher-order EFT terms, referred to as partial-N4LO, to test for their impact on the fitting. The additional terms proved not to be required by the data and did not improve the fit; this result was taken to indicate that corrections from omitted terms are not large compared with experimental errors below 2 MeV. Scattering data were not considered in this work apart from very broad priors on scattering length and effective range; the correlated errors in the S-factor data were taken into account.
Like halo EFT, the phenomenological R-matrix approach avoids a model of nuclear interactions and uses a systematic parameterization to fit data [2010RPPh...73c6301D]. At the time of 2011RvMP...83..195A, had been the subject of very little R-matrix fitting, apparently consisting only of the simple treatment in 2004ADNDT..88..203D that focused on BBN energies. After 2011RvMP...83..195A an R-matrix analysis using the AZURE2 code [2010PhRvC..81d5805A, azure2-manual] was carried out on both elastic scattering and S-factor data in conjunction with the Notre Dame experiment [2013PhRvC..87f5804K], and a more elaborate analysis using Monte Carlo sampling to estimate errors was later reported in 2014PhRvC..90c5804D. In the latter analysis both the fitted value and error bars were heavily influenced by the numerous scattering data of 1964NucPh..50..629B. Very recently, the BRICK software package has been constructed to carry out Bayesian parameter estimation for AZURE2 and applied to both capture and elastic scattering at all energies in the 7Be system by 2022FrP....10.8476O. In this work it was shown that markedly different values result from inclusion or not of the older scattering data [1964NucPh..50..629B] alongside the very recent SONIK scattering data of 2024PhRvC.109a5802P.
A small amount of additional theoretical work in more traditional frameworks has appeared since 2011RvMP...83..195A. This includes fits of potential models to S-factor data [2018PhRvC..97c5802T, 2021NuPhA100622108T] and some calculations of the RGM type [2017PhRvC..96f4605S, 2019PhRvC..99e4618S].
VI.3 Experimental progress on
Since 2011RvMP...83..195A, five new experiments have been reported: Four from the Madrid and ATOMKI groups by the activation technique, at relatively high center of mass energies [2012PhRvC..86c2801C, 2013NuPhA.908....1B, 2019PhRvC..99e5804S, 2023PhRvC.108b5802T], and one from the Notre Dame group using the prompt- method [2013PhRvC..87f5804K].
Following the approach adopted in 2011RvMP...83..195A, we again only use the recoil and activation data for the fits below. The -ray angular distribution is not known experimentally [2007PhRvC..76e5801B], and the resultant uncertainty increases the common-mode error for the prompt- studies somewhat. (An experiment to address previously-raised concerns about angular distribution [2007PhRvC..76e5801B] has recently concluded at Felsenkeller Dresden. The data suggest a higher than expected anisotropy but are so far only available in the form of a PhD thesis [Turkat23-PhD]. If confirmed, they may lead to corrections of a few percent for some of the in-beam experiments where only one angle was instrumented.) In addition, all of the “modern” works reporting prompt- data except for Notre Dame [2013PhRvC..87f5804K] also include data obtained with other methods, leading to partial correlations between data sets that would complicate fitting.
Following this restriction, our data selection proceeds as follows. First, the four data sets previously used in 2011RvMP...83..195A are carried over here: Weizmann [2004PhRvL..93z2503S], LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G] (only the activation data), Seattle [2007PhRvC..76e5801B] (only the activation data), and ERNA [2009PhRvL.102w2502D].
Two of the archival data sets excluded by the 1998 cutoff date in 2011RvMP...83..195A merit some further discussion here: The data of 1982PhRvL..48.1664O, 1984NuPhA.419..115O consist of two points measured by activation in a 3He gas cell. However, their uncertainties are not separated into statistical and systematic components. Including them would have required an uncertain guess on how to divide the error bars, and the fit result would barely change, so we left them out. The work by 1983PhRvC..27...11R consists of one data point that was obtained by averaging two separate activation measurements with a 3He and a 4He gas cell, respectively. We do not use this result because we lack details to verify the background subtraction, an issue that was raised during our meeting by one of its authors. We also choose not to use the activation study by 1983ZPhyA.310...91V. In that experiment there were thick entrance foils and the 4He beam was completely stopped inside a high-pressure 3He gas cell. This gave an integrated measurement over a wide energy range, so that the analysis depends strongly on the assumed shape of the curve, and the result was reported only as an extrapolated .
We now consider the new data since 2011RvMP...83..195A. An activation experiment in Madrid [2012PhRvC..86c2801C] used a 3He beam incident on a 4He gas cell, reporting three data points. Another activation experiment was reported by the ATOMKI group, using a 4He beam on 3He gas cells [2013NuPhA.908....1B]. Two higher-energy campaigns at ATOMKI were again performed using the activation method. These latter data are at = 2.5-4.4 MeV [2019PhRvC..99e5804S] and = 4.3-8.3 MeV [2023PhRvC.108b5802T], respectively, above the energy range suitable for halo EFT, and therefore not included.
A detailed study by the Notre Dame group reported 17 data points [2013PhRvC..87f5804K], using the primary -ray from ground state capture and the secondary -ray from the deexcitation of the 429 keV first excited state of 7Be. No activation data are reported in that work, and the -ray detector was placed at just one angle, 90∘. Since only prompt- data are reported, we did not include this data set. In order to test the effects of this decision, we repeated some of our fits (see below, Section VI.4) with the modern prompt- experiments included: the 17 points from 2013PhRvC..87f5804K, 3 prompt- points from LUNA [2007PhRvC..75f5803C] and 8 from Seattle [2007PhRvC..76e5801B]. The extrapolated S-factor changed by less than 1%, well within the error bars for the recommended .
Finally, an indirect experiment using the 6Li(3He,d)7Be reaction and the ANC technique has recently been reported [2020PhLB..80735606K]. It is left out of our fits due to the additional normalization and theory uncertainties involved in determining an ANC from a transfer experiment, which are larger than the errors in the S-factor measurements used for the present fits. All the data used in the present fits are summarized in Table 5.
As in 2011RvMP...83..195A, we model data uncertainties as consisting of a component that is independent for each point and a common-mode component that applies to all data from a given experiment as a multiplicative factor. This separation is well-documented for all of the modern data and is also shown in Table 5. Except at the lowest energies, the common-mode error typically dominates.
| Group and references | Energy | Data points | Common | Rescaling factor | ||||
| range [keV] | used | total | mode uncert. | Higa EFT | Zhang EFT | BRICK | BRICK+S | |
| Weizmann [2004PhRvL..93z2503S] | 4 | 4 | 1.03(2) | 1.02(2) | 1.02(2) | 1.03(2) | ||
| LUNA [2006PhRvL..97l2502B, 2007PhRvC..75f5803C, 2007PhRvC..75c5805G] | 7 | 7 | 1.02(2) | 1.04(2) | 1.05(2) | 1.01(2) | ||
| Seattle [2007PhRvC..76e5801B] | 8 | 8 | 0.96(1) | 0.95(2) | 0.95(2) | 0.98(2) | ||
| ERNA [2009PhRvL.102w2502D] | 27 | 27 | 0.96(3) | 0.94(2) | 0.96(2) | 0.99(2) | ||
| Madrid [2012PhRvC..86c2801C] | 2 | 3 | 0.99(3) | 0.99(3) | 0.97(2) | 0.99(2) | ||
| ATOMKI [2013NuPhA.908....1B] | 4 | 5 | 1.01(3) | 1.00(3) | 1.01(3) | 1.04(3) | ||
VI.4 Data fitting
We base our recommended on fits to the halo EFT and R-matrix parameterizations discussed above. These avoid tacit assumptions present in potential-models and conceptual difficulties involved in combining ab initio constraints with data. The fits presented here differ from the previously published fits of 2020EPJA...56..166P, 2020JPhG...47e4002Z, and 2022FrP....10.8476O mainly in the uniform use of the agreed-upon capture data and uncertainties from Section VI.3 across all fits. We restricted fitting to the MeV range of validity for the NLO halo EFT expressions. Despite uniform handling of capture data, the fits in each framework handle scattering inputs differently for reasons discussed below.
In addition to the total capture cross section and scattering data, the fitted data also include branching ratios for capture into the two 7Be bound states, taken from 2007PhRvC..76e5801B, 2009PhRvL.102w2502D, 2007PhRvC..75f5803C, and 2013PhRvC..87f5804K. These are necessarily from prompt- experiments and suffer from the concerns about angular distribution discussed in Section VI.3. However, their inclusion simplifies the fitting considerably by breaking parameter degeneracies between ground- and excited-state transitions (especially in the fitted ANCs), probably without strong impact on .
In constructing fits we split common-mode and point-to-point errors and “float the norms” of data sets using the cost function
to describe goodness of fit [1994NIMPA.346..306D].
A rescaling factor is fitted for each data set , with its deviation from unity penalized by the common-mode errors given in Table 5. The index sums over all points within a given data set; is a measured cross section, is a predicted cross section, and is the point-to-point error of the th point in data set .
We carried out both frequentist fits that minimize and Bayesian fits based on the posterior probability distribution of a likelihood function computed from . In the Bayesian fits parameters, extrapolated , and their errors were determined by Monte Carlo sampling of the parameter space. The relatively large number of model parameters and the significant fitting degeneracies between some of them make Bayesian analysis a natural choice for finding best fit and confidence intervals for the multi-parameter models applied here. Two of the groups ran into serious difficulty with the frequentist fits due to parameter degeneracies and shallow local minima; during this study they were unable to produce frequentist fits in which they had confidence.
We performed multiple fits in the NLO halo EFT of 2018EPJA...54...89H and 2020EPJA...56..166P. These fits were carried out both with and without scattering constraints, which mainly impact by removing parameter degeneracies that would otherwise leave the -wave scattering length poorly constrained. The code base for this version of halo EFT incorporates scattering data through phase shifts. These were taken from the partial-wave analysis of 1972NuPhA.195..241B, which at low energy are based mainly on the data of 1964NucPh..50..629B.
One set of fits for the 2018EPJA...54...89H halo EFT was produced by minimization, proceeding in two steps: first LO parameters were fitted to keV data, and then the results were taken as initial values in the search to minimize over all parameters in the NLO theory for the full set of data. These fits gave keV b with the scattering constraint and keV b without (uncertainties being propagated in linear approximation using covariances and partial derivatives). Formally the EFT truncation error from stopping at NLO corresponds to an additional theoretical error that can shift S-factors by 10%. However, any fitted curve is constrained by low-energy data, which in some sense become effectively renormalization conditions of the field theory; the error on extrapolated should probably be smaller than 10% by an amount that is hard to estimate.
A second set of fits to the 2018EPJA...54...89H halo EFT was carried out using Bayesian methods. Priors for the EFT parameters were developed based on previous experience, and data rescaling factors were incorporated as additional priors. The Bayesian results (including only experimental error) are keV b including the phase shifts by 1972NuPhA.195..241B and keV b excluding them, consistent with the -minimization.
Searches of the parameter space to minimize for the NLO halo EFT of 2020JPhG...47e4002Z ran into difficulties with parameter degeneracy and local minima in the surface; for this formalism we report only Bayesian results. No experimental information about scattering was used, but flat priors on scattering length and effective range were chosen over a 5 range around recent experiment. This fit differs from 2020JPhG...47e4002Z mainly by excluding the Notre Dame data and including the ERNA activation data, and it gives keV b. The error associated with EFT truncation at NLO is estimated in this approach by separately fitting the partial-N4LO theory discussed above. The result suggests that EFT truncation at NLO affects extrapolation from the data to threshold by –3%.
Our R-matrix fits are based mainly on sampling Bayesian posterior probabilities with the BRICK code. We also produced frequentist fits, but we were unable to estimate their errors convincingly. The R-matrix fits used capture data both alone and in combination with the SONIK scattering data [2024PhRvC.109a5802P] – the latter being chosen because of concerns with the 1964NucPh..50..629B data that are discussed in 2022FrP....10.8476O. Since BRICK fits elastic differential cross sections directly, it was not feasible to use the same phase-shift-based scattering constraints as our halo EFT fits. The R-matrix fit that includes scattering data is essentially the “CS” fit of 2022FrP....10.8476O, but with a restriction to only the capture data described in Table 5, only scattering data below 2 MeV center-of-mass, and only the R-matrix parameters relevant below 2 MeV (no level or radiative widths for -wave background poles).
| Fit | ||||
| [keV b] | [MeV-1] | [fm] | [fm-1] | |
| 2018EPJA...54...89H EFT-NLO | 0.554 | -0.56 | 30 | |
| 2020JPhG...47e4002Z EFT-NLO | 0.573 | -0.61 | 46(6) | 29(3) |
| BRICK R-matrix | 0.562 | -0.57 | 24.4(9) | |
| BRICK R-matrix+scattering | 0.531 | -0.47 | 22.6(7) | |
| ANC [2020PhLB..80735606K] |
Several of the fits are compared in Fig. 7. There it is apparent that they all agree within 5% in the 0–2000 keV energy range, and within 4% in the astrophysically relevant range 0–500 keV. The fit to the 2018EPJA...54...89H NLO EFT lies in the middle of the range, so we adopt it as a reference in the lower panel of the figure. Relative to that fit, the fits to the NLO EFT of 2020JPhG...47e4002Z and the R-matrix fit without scattering rise up as much as 4% higher at 0–200 keV.
The most striking difference among the curves is between the BRICK R-matrix fit including the SONIK scattering data [2024PhRvC.109a5802P] and all the other fits. While the R-matrix fit with scattering finds rescaling factors somewhat closer to unity than the other fits do, its energy dependence in the 0–1000 keV range is qualitatively different, as is especially visible in the lower panel of Fig. 7 where residuals relative to the fitted 2018EPJA...54...89H EFT are shown. The BRICK fit without the scattering data is much closer to all the halo EFT fits, so the recent scattering data apparently have a large impact.
VI.5 Recommended value
Because the fit to the NLO EFT of 2018EPJA...54...89H gives the central result among those attempted here, we adopt it for our recommended . For ease of adoption, it is noted that this recommended curve can be empirically parameterized using the same shape as in 2011RvMP...83..195A, by the numerical equation (following a customary form in the past literature [1987ApJ...319..531K, 2011RvMP...83..195A])
Here is the center of mass energy in MeV. Equation (VI.5) reproduces our recommended curve within 0.3% and is only applicable for = 0–1600 keV.
Propagation of uncertainties from measurement to extrapolated is relatively straightforward and unambiguous (at least in the Bayesian fitting). However, the uncertainty due to theoretical approximation is more complicated, and error estimation on extrapolated quantities in EFTs remains an open area of research; estimates from the groups working on are discussed above. We estimate the overall theoretical uncertainty on our recommended S-factor to be the measured by the 4% dispersion among fitted curves; this is comparable to the more formal estimates.
Our recommendations for the zero energy astrophysical S-factor and the low-energy slope are then
The given value of applies over = 0-500 keV. Its considerable uncertainty is given by the full span of slopes among the four fits discussed above, which is dominated by the inclusion of elastic scattering in one of the R-matrix fits. The curve is quite straight at low energy in all models; its second derivative is consistent with zero and even difficult to compute for a given fit because of numerical cancellations.
The first error bar for is given by the fit uncertainty, conservatively using the largest value among the various fits, and it encodes the experimental uncertainty. The second error bar reflects the 4% dispersion among theoretical formalisms. The two errors can be added in quadrature for applications. Our recommendation is essentially the same as in 2011RvMP...83..195A, but derived from completely different theoretical approaches that are much more directly connected to the data and much less dependent on model assumptions. It is also based on data over a wider energy range. The dispersion of and among models at the solar Gamow peak is illustrated in Table 6, along with some indication of how fitted scattering lengths and squared ANCs vary among frameworks. (The latter are highly degenerate with short-range parameters in the fitting; in the 2018EPJA...54...89H EFT they are sensitive to small variations in the -wave effective ranges that are fitted instead of ANCs.)
VI.6 Recommendations for future work on
Here we present recommendations for further work that would improve understanding of the 3He()7Be reaction. Our first recommendation for experiment concerns elastic 3He+4He scattering. New experiments over a wide energy range including at least part of the astrophysical range are needed to address the marked impact of the recent SONIK scattering data [2024PhRvC.109a5802P] on fitting, for independent confirmation. Data below 2 MeV would be most useful for consistent fitting of halo EFTs.
Second, new measurements (besides the preliminary data of Turkat23-PhD) are needed for the angular distribution of the emitted -rays. This would relieve concerns about including prompt- data, most notably the Notre Dame [2013PhRvC..87f5804K] that do not include activation measurements, in future fits. These experiments should also be designed to provide new angle-corrected branching ratios for capture into the two bound states.
Our third recommendation is a study of the astrophysical S-factor over a wide energy range, 50-1500 keV and extending to even lower energies when technically feasible, with small point-to-point errors. Even with a large common-mode error, this would significantly constrain theoretical curve shapes near threshold, and providing an independent check on the influence of scattering data on .
On the theoretical side, further development of ab initio reaction calculations is desirable, improving on points of current difficulty like explicit inclusion of three-body potentials. Calculations with the same potential but by different methods would help to separate computational issues from the impact of potential choice. Further exploration of how ab initio inputs like ANCs and scattering lengths could serve as fitting priors would also be a useful avenue of research.
Authors of future theoretical calculations should consider including -ray angular distributions in their results; until very recently angular distributions were only available from two potential models [1981PhRvC..23...33K, 1963PhRv..131.2582T]. New calculations would inform analysis of experiments, and the angular distribution measurements we recommend would provide tests of the calculations (or fitting constraints).
A source of complication here has been the need for a good meeting point between scattering data and theory. Given the difficulties of constructing a unique phase shift analysis without an underlying theory, and the recent arrival of modern scattering data, theorists should consider including elastic differential cross sections directly as outputs and fitting constraints.
The hep reaction ()
The reaction , also known as the hep reaction, is one of the possible processes taking place after the , and it produces the most energetic neutrinos, with an endpoint energy of 18.8 MeV. It is however the least probable, with its S-factor, and consequently its rate, even beyond the reach of current experiments. This is due to the fact that the hep reaction is induced by the weak interaction and further suppressed by Coulomb barrier and by the fact that the leading one-body axial (Gamow-Teller) current operator cannot connect the main -state components of the and initial- and final-state wave functions. As a consequence, the one-body axial current non-zero contribution is due to the small components of the and wave functions. Furthermore, other contributions, such as weak magnetism and other one-body corrections to the vector current, and two-body vector and axial currents, and the -waves in the initial state, normally suppressed at solar temperatures, are significantly enhanced. This is further complicated by cancellations between the one-body and two-body axial current contributions. As a consequence, similarly to the pp fusion, the hep reaction cross section is too small to be measured in laboratory and only theoretical predictions are available.
The most recent studies of this reaction, already evaluated in 2011RvMP...83..195A, are those of 2000PhRvC..63a5801M and 2003PhRvC..67e5206P. The first one was performed within a phenomenological approach, similar to the calculation of 1998PhRvC..58.1263S for the pp fusion. In this particular case, the and initial and final nuclear wave functions were obtained with the correlated Hyperspherical Harmonics (CHH) variational method [1995FBS....18...25V, 1998PhRvL..81.1580V], using the Argonne (AV18) two-nucleon potential [1995PhRvC..51...38W] augmented by the Urbana IX (UIX) three-nucleon interaction [1995PhRvL..74.4396P]. In this way, the binding energies of 3He and 4He, and the singlet and triplet scattering lengths were in good agreement with experiment. The weak vector and axial transition operators were obtained within a phenomonological approach. In particular, the largest two-body contribution to the axial current arising from the excitation of intermediate -isobar degrees of freedom was included using the transition-correlation operator scheme of 1991PhRvC..44..619C, 1992PhRvC..45.2628S and fixing the nucleon-to- axial coupling constant to reproduce as in the pp case [1998PhRvC..58.1263S].
The second study of 2003PhRvC..67e5206P was performed within the same hybrid EFT approach used for the pp funsion (EFT∗), with the same nuclear weak current operators obtained within EFT as discussed in Section III, the LEC being fixed to reproduce . The initial and final wave functions were derived using the CHH method with the AV18/UIX potential, as in 2000PhRvC..63a5801M.
Following the results of 2000PhRvC..63a5801M and 2003PhRvC..67e5206P, 2011RvMP...83..195A recommended for the zero-energy S-factor, , the value keV b. Three recommendations were also made: (i) to perform new studies with a broad spectrum of Hamiltonian models, in order to properly access the theoretical uncertainty; (ii) to study other weak reactions for which experiments can test the predictions made by employing the same theoretical ingredients as those used for the hep reaction calculation; (iii) to further understand the relation between the hep reaction and the so-called hen process (). Of these recommended works, only the second one has been performed so far, applying the theoretical framework used for the hep study to the muon capture reactions on light nuclei [2011PhRvC..83a4002M].
Given the lack of further studies on the hep reaction after 2011RvMP...83..195A, we decide to maintain
as recommended value. Notice that we could have increased by %, in analogy with , due to the updated values of input parameters (mostly ). However, we have decided not to apply such an increase, because it does not derive from a direct calculation, and, most of all, because such an increase lies well within the quoted % uncertainty. However, given the advances in ab initio studies of the last decade and in the development of the EFT framework, we update the recommendations of 2011RvMP...83..195A as follows: (i) the hep process should be revisited within a fully consistent EFT approach, similarly to what has been done in these years for the pp fusion and other processes, as, for instance, muon captures on light nuclei [2012PhRvL.108e2502M, 2018PhRvL.121d9901M, 2018PhRvC..98f5506A, 2023FrP....1049919C, 2023PhRvC.107f5502B, PhysRevC.109.035502]; (ii) the initial and final wave functions should be calculated with the most accurate ab initio methods, as the (uncorrelated) HH method [2008JPhG...35f3101K, 2020FrP.....8...69M], which has been proven to provide more accurate bound and scattering wave functions than the CHH method, using both local and non-local interactions; (iii) the relation between the hep and the hen process should be understood. The hen process has been studied within EFT and the HH method by 2022PhRvC.105a4001V, but a consistent parallel study of both reactions is still missing.
A relatively recent SNO Collaboration analysis of data from all three of the detector’s running phases yielded a one-sided confidence-interval bound on the hep neutrino flux of 2020PhRvD.102f2006A
a result in agreement with the SSM prediction. The lack of a definite measurement, however, has meant that there is no substantive experimental test of the predicted Various global analyses of solar neutrino data have provided weak evidence for a nonzero flux, e.g., from Bergstrom:2016cbh
which is consistent with the experiment bound given above.
An extraction of the hep flux from experiment will likely require a detailed shape analysis to separate hep neutrinos from the high-energy tail of the 8B spectrum and the low-energy tail of atmospheric neutrinos. Typically the hep spectrum is assumed to have an allowed shape: indeed, in the work reported here spectra are not provided, where -waves, weak magnetism, and similar corrections have been included. Because these corrections are unusually large for hep neutrinos, future studies should provide spectra and evaluate the impact of such corrections on the shape.
Electron capture by p+p and
Electron capture reactions are the sources of lines found in the solar neutrino spectrum. In particular, electron capture on , i.e., the process d, also known as the pep reaction, competes with fusion and depends on the same nuclear matrix element. Therefore, the ratio between the pp and pep rates is independent of nuclear physics. Based on this consideration, in 2011RvMP...83..195A the result of 1998RvMP...70.1265A, which is based on the work of 1969ApJ...155..501B, was multiplied by the radiative corrections calculated by 2003PhRvC..67c5502K, leading to the final result given by
where is the density in units of g/cm3, is the mean molecular weight per free electron, and is the temperature in units of MK. The range of validity is . Given that no new evaluation of the pep rate has been performed since 2011RvMP...83..195A, the result of Eq. (20) represents also the present recommended value.
Competition between electron and proton capture on 7Be fixes the branching ratio of the ppII and ppIII chains, and thereby the 7Be and 8B neutrino fluxes. Superallowed electron capture on the ground state of 7Be leads to the ground state or the , 478 keV first excited state of 7Li with a measured (terrestrial) branching ratio of 10.44(4)% [2002NuPhA.708....3T]. It is customary to calculate the solar decay rate in terms of the measured terrestrial decay rate, which has the advantage of removing the nuclear physics dependence, but it requires a proper calculation of the densities of electrons at the nucleus. This has usually been done by considering continuum and bound electrons separately, using the Debye-Hückel (DH) approximation pioneered by 1954AuJPh...7..373S for the screening of bound electrons in the solar plasma; the capture rate of continuum electrons is not significantly influenced by electron screening. Furthermore, it is necessary to calculate the atomic probability densities governing the and terrestrial electron capture rates. 2020PhRvL.125c2701F recently measured the ratio of to capture using superconducting tunnel junctions. The recommendation of both 1998RvMP...70.1265A and 2011RvMP...83..195A was
which is valid for . This expression is based on the continuum capture rate calculated by 1969ApJ...155..511B, the average ratio of the total capture rate to the continuum rate calculated using three solar models in 1994PhRvD..49.3923B, and the current value of the half-life, 53.22(6) d [2002NuPhA.708....3T]. It agrees within 1% with a density matrix calculation of the rate that makes no assumptions regarding the nature of electronic quantum states in the solar plasma and allows for aspherical fluctuations in the spatial distributions of plasma ions [1997ApJ...490..437G]. The estimated uncertainty of 2% accounts for possible corrections to the Debye-Hückel approximation due to thermal fluctuations in the small number of ions in the Debye sphere, and breakdowns in the adiabatic approximation [1992ApJ...392..320J].
Although calculations have appeared in the literature claiming new plasma effects based on model calculations (see 2013ApJ...764..118S, 2019A&A...623A.126V), we continue to regard the arguments in 2002A&A...383..291B as definitive. That reference presented five distinct derivations demonstrating that the Salpeter formula for screening corrections is valid for pp-chain reactions and solar conditions, up to corrections on the order of a few percent.
The reaction ()
IX.1 Introduction
Radiative proton capture on (=3/2-) at solar energies proceeds through non-resonant capture to the ground state of (=2+). This capture occurs predominantly at separations well beyond the range of the strong interaction via transitions from and partial waves, as the amplitude and contributions from higher partial waves are negligible in the energy range of interest [1961NucPh..24...89C]. The importance of this reaction for the determination of the high energy solar neutrino spectrum and the experimental data reviewed by 1998RvMP...70.1265A inspired several experiments that, taken together, provided a consistent picture of the energy dependence of the reaction cross section, despite discrepancies in absolute scale. Prior to the review of 2011RvMP...83..195A, most direct measurements were performed using a radioactive target and an intense proton beam. In spite of the great care taken in evaluating the systematic energy-independent uncertainties, in the following referred to as common-mode errors (CMEs), a remarkable discrepancy persisted even in the most recent experiments, that was finally handled in 2011RvMP...83..195A by inflating the stated experimental uncertainties. Indirect measurements of the cross section of B based on Coulomb dissociation were considered, but not included in the final recommendation of 2011RvMP...83..195A for the zero-energy astrophysical S-factor . However, despite the experimental issues, the uncertainty on the recommended value of in 2011RvMP...83..195A is dominated by the theoretical contribution. Therefore, our analysis focused on these aspects (discrepancy of experimental data, indirect measurements, theoretical uncertainties), as discussed in the next sections.
IX.2 Experimental Data
IX.2.1 Direct Measurements
The thorough review of published work done in 2011RvMP...83..195A investigated the influence of beam-target overlap, target stoichiometry, beam energy loss, and the backscattering of 8B recoils on CMEs and led to the selection of a homogeneous group of well-documented data sets. Since then, no further information became available on these aspects of direct measurements. The data that formed the basis of the 2011RvMP...83..195A recommendation were those of 2003PhRvC..67f5805B, 2003PhRvL..90b2501B, 1998PhRvL..80..928H, 2001PhRvL..86.3985H, 1999PhLB..462..237H, 2001NuPhA.696..219S, and the BE3 data set of 2003PhRvC..68f5803J. We considered the complete set of radiative capture measurements that were the basis of the 2011RvMP...83..195A recommendation for , where we adopted the revised BE3 data presented by 2010PhRvC..81a2801J, which were reanalyzed using a more sophisticated model of the target and more accurate stopping powers. In addition, we included in our analysis the only new experiment reported thereafter of 2022PhLB..82436819B, which used a radioactive ion beam and the recoil mass separator ERNA to detect the 8B recoils. While these new results contribute to the determination of with different, well controlled systematics [2018EPJA...54...92B], their large statistical uncertainty relative to the other experiments, due to the low counting statistics, limits their impact.
New scattering data have been published. Elastic and inelastic scattering cross sections at relative kinetic energies between 474 keV and 2.74 MeV were measured and wave scattering lengths were inferred by 2019PhRvC..99d5807P. At higher energies between 1.6 and 3.4 MeV, thick target elastic and inelastic scattering excitation function measurements were used to infer the existence of new resonances above 1.8 MeV [2013PhRvC..87e4617M]. These high lying resonances have no direct influence on because of their small partial widths in the channel and/or because they cannot be populated by s- or d-wave capture at low energy.
IX.2.2 Indirect Measurements
In 2011RvMP...83..195A, Coulomb dissociation measurements based on the formalism developed by 1986NuPhA.458..188B were considered [PhysRevLett.83.2910, PhysRevLett.73.2680, KIKUCHI1997261, 1998EPJA....3..213K, PhysRevLett.86.2750, PhysRevC.63.065806, PhysRevLett.90.232501, 2006PhRvC..73a5806S], where in some cases the most recent publications supersede the previously published ones. Finally, the results of these experiments were not included in the determination of due to disagreements over whether the analyses of the various measurements properly accounted for the contributions of transitions. We disagree with these reservations and find the agreement of these measurements with radiative capture measurements significant. However, in order to include these consistent data in formulating our recommendation for , we inflated the CMEs of the Coulomb breakup measurements to adequately account for the uncertainty in the E2 components assumed in each analysis.
The common mode errors are obtained from a linear sum of those given in the manuscripts and an additional common mode error due to the estimated size of the component in each measurement. This additional error is applicable in one direction only, downward for the 1998EPJA....3..213K and 2006PhRvC..73a5806S analyses of the RIKEN and GSI measurements, respectively, and upward for the 2003PhRvC..68d5802D analysis of the Michigan State University (MSU) measurement, according to whether the component was included or excluded. Though the 2003PhRvC..68d5802D analysis did include a common mode error contribution of 2.5% due to the component (which represented 5% of the total cross section), we have included an additional 2.5% contribution to the MSU experiment. The size of the additional common mode errors for the RIKEN and GSI measurements was based on scaling this 2.5% according to the predicted size of the contribution to the cross section in these measurements relative to that in the MSU measurement, evaluated using first-order perturbation theory and the and S-factors calculated in a potential model of 8B [1996ZPhyA.356..293B]. Thus the additional single-sided common mode errors added to the GSI, MSU, and RIKEN measurements are -2.1%, 2.5%, and -4.3%, respectively.
IX.3 Theory
The size of the theoretical contribution to the uncertainty in the 2011RvMP...83..195A estimate of was twice that of the experimental contribution. Since then, significant progress has been made. A new potential model calculation was performed by 2019NuPhA.983..175D. The large-scale computational demands of calculations using high-quality nucleon-nucleon (NN) and three-nucleon (3N) interactions beyond =5 had hindered the ab initio approach in the past. Nevertheless progress has been rapid and it has been possible to extend ab initio calculations to B, albeit without 3N forces. 2011PhLB..704..379N performed such a calculation using the no-core shell model in the continuum with a nucleon-nucleon interaction derived from chiral EFT at fourth order. The calculation converged with respect to the model space size and reproduces experimental data for the S-factor despite having virtually no free parameters (though the similarity renormalization group evolution parameter was tuned to reproduce the binding energy of 8B). The final value of =19.4(7) eV b agrees with the 2011RvMP...83..195A recommendation and, as will be shown later, with our analysis. More recently, a new approach was introduced to combine experimental measurements with ab initio predictions, resulting in an “ab initio-informed evaluation” Kravvaris23-PLB that arrived at a value of eV b.
A significant theoretical advance was the application of the formalism of halo EFT developed by 2002NuPhA.712...37B to 7Be(B calculations by 2015PhLB..751..535Z, 2018PhRvC..98c4616Z, and 2022PhRvC.106a4601H. It treats the incoming and 7Be as point-like particles and the 8B final state as a bound state of the two. The calculation and the associated errors depend upon the accuracy with which one is able to describe the incoming scattering states, the final bound state and the relevant electromagnetic currents. The relative kinetic or CM energy range MeV, corresponding to a relative momentum MeV/c, is considered within the domain of applicability of halo EFT. The two cluster physical description of is expected to hold at 3He- relative momenta below a physical cutoff MeV/c set by the threshold (binding momentum) for breaking the 7Be core into these constituents. Further, at energies above the excitation energy of the first excited state, MeV, including the contribution of the excited core is imperative in the EFT.
2015PhLB..751..535Z and 2018PhRvC..98c4616Z include such an excited contribution. 2022PhRvC.106a4601H provides a calculation without the contribution below and one with it above that also includes an contribution to the resonance. The two groups’ halo EFT calculations with the contributions differ in several respects that are elaborated in 2022PhRvC.106a4601H. Differences in the final bound state calculation affect the treatment of the short-range interactions and divergences in the EFTs. The interpretation of the asymptotic normalization constants (ANCs) in terms of elastic scattering parameters is affected. However, the momentum dependence of the capture cross section is not impacted by the ANCs, which are fitted to capture data. Thus there is no effect on the cross section. In the incoming channel, 2022PhRvC.106a4601H use a more general form of the -wave short-ranged interaction involving the core based on the low-energy symmetry that is included only at NNLO. In contrast, 2015PhLB..751..535Z and 2018PhRvC..98c4616Z include a short-ranged interaction at LO that could make a difference in the momentum dependence of the cross section. However, as shown in an order-by-order calculation [2022PhRvC.106a4601H], the short-ranged interaction is a sub-leading effect so the two halo EFTs should have similar low-momentum dependence. The third important difference is in the treatment of two-body currents, which affects EFT error estimates. 2015PhLB..751..535Z and 2018PhRvC..98c4616Z include this at NLO whereas 2022PhRvC.106a4601H estimate it to be an N3LO effect. Bayesian estimates of the two-body current [2015PhLB..751..535Z, 2018PhRvC..98c4616Z] are not strongly constrained by capture data, consistent with this being a higher-order contribution.
Another widely used approach for describing the cross sections of low-energy nuclear reactions is R-matrix theory. In R-matrix theory the 7Be(B cross section can be parametrized in terms of a few parameters representing either “real” or “background” poles. The real poles correspond to the observed states and resonances, while the background terms correspond to the “mean-field” effects that are of the non-resonant type. In the present case for example, a background pole in the 1- channel is needed so as to reproduce the non-resonant part of the capture. The parameters of the real pole corresponding to the bound state are determined by the binding energy and the two ANCs, while the 1+ resonance parameters are determined by its energy and partial decay widths (both proton and radiative). The R-matrix analysis of this reaction carried out by 1995NuPhA.588..693B arrived at a value of eV b; however, it was based on data that were excluded in 2011RvMP...83..195A and here, with the exception of those of 1983PhRvC..28.2222F.
We fit data using R-matrix theory and halo EFTs, with both frequentist and Bayesian approaches, investigating the impact of different choices for the energy range in which data were fitted, and obtained consistent results. Our recommended value and its uncertainty are based on a Bayesian fit of both direct and Coulomb dissociation data taken at energies using the 2022PhRvC.106a4601H halo EFT⋆. This theory provides a good description of capture data over a wide energy range including the contribution from the resonance. The theoretical uncertainties are well understood. Further, the contributions of the and spin channels to the capture cross section are self-consistently parametrized with the elastic scattering information contained in the known ANCs for this system.
The Bayesian formalism naturally incorporates prior information concerning the ANCs in the fits. The upper limit on the energy in the fit was determined to be high enough to include some data from all the modern experiments, and yet low enough that the contribution of the wide resonance at does not exceed of the total cross section at any energy, as estimated by our R-matrix analysis. The fitting procedure is described in the next section, whereas the details, as well as the results of the fits done in different energy ranges and with different theories are reported in the supplemental material.
IX.4 Fitting Procedures
In 2011RvMP...83..195A, it was noted that the different data sets had similar energy dependences. The discrepancies among the data sets were primarily due to different absolute normalizations. We address the normalization issue by introducing scaling factors associated with the CMEs. This contrasts with the 2011RvMP...83..195A approach, in which the errors were inflated by the factor . The procedure for handling discrepant data sets adopted here is mathematically rigorous and aligns naturally with the method of estimating theoretical uncertainty discussed below. We prefer this approach to the inflation factor method used in 2011RvMP...83..195A, since the absolute scales of the different data sets can be used as constraints for each other, while there is no information available for a reevaluation of the CMEs. The procedure we used, described below, can be applied to both frequentist and Bayesian analyses.
Although our final recommendation is based on Bayesian analysis alone, both minimization (a frequentist approach) and Bayesian posterior probability distribution function evaluation were performed to fit EFTs to the data; in the R-matrix analysis, only the former was used. The likelihood function, defined as the conditional probability for the data given theoretical parameters , enters both fitting procedures. It can be derived from the expression given in 1994NIMPA.346..306D
with labeling the different experimental data sets and labeling the individual data points of each experiment. In the second term, we divide theoretical predictions for the data points of the experiment by a scaling factor and consider the combination as our full model for , the th data point of experiment with point-to-point uncertainty .
In the first term, we assign a Gaussian distribution centered at with a width equal to the CME ; for the GSI, MSU and RIKEN data shown in Table LABEL:table:normsBayesian, asymmetric Gaussians are used with depending on the sign of . In our minimization, the definition is derived from equating the full exponents in Eq. 22 to . In the absence of CMEs with , Eq. 22 leads to the usual uncorrelated least-squares minimization formula.
In the Bayesian analysis, the terms in the first exponential are considered as part of the prior distribution that are conditioned on prior knowledge and assumptions such as the CMEs in the scaling factors or estimates of theoretical parameters in the EFT. The posterior distribution for the theoretical parameters is then derived via Bayes’s theorem,
where the likelihood function is defined without the first exponential in Eq. (22). However, since the GSI, MSU and RIKEN data have asymmetric CMEs (see Table LABEL:table:normsBayesian), we draw priors from a uniform distribution spanning 50% to 150% of the central values for these that are then used in the asymmetric normalization exponentials in the likelihood function. This was computationally simpler than generating asymmetric Gaussian priors that span the entire range, which is neither necessary nor physical for a normalization constant.
The constant in Eq. (23), known as the Bayesian evidence, guarantees the correct normalization of the posterior distribution. It is useful in comparisons of different theories but does not affect the estimation of parameters.
IX.5 Data Analysis and Determination
We performed fits using several theoretical expressions over multiple energy ranges: , , and over a “non-resonant” energy range and , in which the contribution from the resonance can be ignored. Fits with and without Coulomb dissociation data were performed that resulted in overlapping determinations within the estimated fitting uncertainties. The different low energy fits included one using the halo EFT from 2018PhRvC..98c4616Z and one using the halo EFT⋆ from 2022PhRvC.106a4601H, both including the excited contribution. The simpler halo EFT expressions, not including an excited component , were also fitted in this low energy region with comparable results. For the fits in the other two energy ranges, halo EFT expressions with excited contributions from both 2018PhRvC..98c4616Z and 2022PhRvC.106a4601H were used. The 2018PhRvC..98c4616Z halo EFT expression, which does not include an contribution, was supplemented with a Breit-Wigner resonance [1988ccna.book.....R] with its energy (around the resonance energy ), width, and proton-decay branching ratio determined by the fits. An R-matrix model was fitted to data in the three energy regions with the AZURE2 code [2010PhRvC..81d5805A], again resulting in an consistent with the halo EFT values. The results from the various fits are included in the supplemental material.
Bayesian fits were performed with halo EFT expressions from both 2020JPhG...47e4002Z and 2018EPJA...54...89H that are compatible with the various fits. The Bayesian analysis had advantages over the fits in exploring parameter space and in determining realistic uncertainties [PhysRevLett.122.232502].
The capture proceeds through the and channels. The reaction is known to be peripheral, resulting in a that has a sub-leading dependence on the strong interaction in the initial state [2000PhRvC..62f5803B]. This behavior was shown to persist away from the threshold in halo EFT, in which for capture the strong interaction in the incoming -wave channel only contributes at NNLO.
Therefore, to high order, the capture is only sensitive to the sum of the squares of the ANCs in the and channels. This results in an arbitrariness in the relative contributions from the two spin channels in the fits, without affecting the final determination. R-matrix, EFT, and EFT⋆ fits confirm this behaviour. On the other hand, the ANCs have been extracted experimentally [2003PhRvC..67f2801T, 2006PhRvC..73b5808T] and also theoretically [2011PhRvC..83d1001N, 2018PhRvC..98c4616Z] in ab initio calculations. Although the determination of is insensitive to the relative contributions of the two spin channels, knowledge of the relative contributions of the ANCs is important in the EFT framework. It establishes the hierarchy of different contributions in the perturbative expansion. This is crucial in developing a self-consistent expansion that is necessary to quantify the theoretical errors at any finite order of the perturbation. Prior knowledge of the ANCs and other parameters can be accounted for naturally in the Bayesian framework in which one specifies the prior probabilities . Similar constraints can be included in the fit by modifying the first exponential in Eq. (22) to center the theoretical parameters around known values.
The Bayesian fits involved six theoretical parameters and 11 normalization constants for the data sets. The priors for the normalization constants were as described earlier. The priors for the theoretical parameters are based on the underlying EFT assumptions [2022PhRvC.106a4601H] about the expansion in a small ratio where represents the low momentum of infrared physics and the cutoff of the theory. The and scattering lengths fm and fm [2019PhRvC..99d5807P] are taken as inputs. For capture at LO, the only fit parameter, the effective momentum is assumed to scale as and accordingly we assume a uniform prior for between -100 MeV/ and 1.5 MeV/, where the upper limit is set by the physical constraint that the ANC squares have to be positive. At NLO, the ANC2 is a fit parameter whose prior is taken from a normal distribution determined by experiment. At NNLO, we use the ab initio calculation of [2018PhRvC..98c4616Z] to draw the ANC2 from a normal distribution. A ratio of the -wave scattering lengths in the coupled-channel - that is assumed to be in the EFT is drawn from a normal distribution with a mean of 0 and a standard deviation of 10 to cover a wide range. The capture is dominated by the transition and it requires three parameters: a -wave scattering volume , an effective momentum and a two-body current to regulate divergences. 2022PhRvC.106a4601H determined and from the the narrow resonance energy and width. Here, we kept fixed while drawing from a uniform distribution between -150 and -50 MeV. The Bayesian fit gave . The two-body coupling was drawn from a uniform distribution and the fits gave a numerical value consistent with the EFT estimate. We obtain the posterior distribution using a probabilistic integration method called Nested Sampling [Skilling06-BA] implemented in Python [2009MNRAS.398.1601F] that calculates both the posterior and the evidence.
The experimental data and EFT⋆ fits at LO, NLO, and NNLO are shown in Fig. 8. and its derivatives at threshold are given in Table 7. The quality of the fits, the consistency of the numerical values of the fit parameters with EFT assumptions, and the order-by-order improvement in the numerical value of the fitted capture cross section and S-factor give us confidence in the estimated theoretical uncertainty. We find the posterior probability distribution functions to be symmetric and well described by Gaussians, and as such we report the mean and standard deviation in Table 7. Table LABEL:table:normsBayesian gives the scaling factors for the experimental data sets determined from the NNLO fit along with their common mode errors.
| Theory | () | () | () |
[…]
| Data set | Scaling | Data points | CME |
[…]